### v1

# Can Uncertainty Quantification Enable Better Learning-based Index Tuning?

Abstract
Index tuning is crucial for optimizing database performance by selecting optimal indexes based on workload. The key to this process lies in an accurate and efficient benefit estimator. Traditional methods relying on what-if tools often suffer from inefficiency and inaccuracy. In contrast, learning-based models provide a promising alternative but face challenges such as instability, lack of interpretability, and complex management. To overcome these limitations, we adopt a novel approach: quantifying the uncertainty in learning-based models’ results, thereby combining the strengths of both traditional and learning-based methods for reliable index tuning. We propose Beauty, the first uncertainty-aware framework that enhances learning-based models with uncertainty quantification and uses what-if tools as a complementary mechanism to improve reliability and reduce management complexity. Specifically, we introduce a novel method that combines AutoEncoder and Monte Carlo Dropout to jointly quantify uncertainty, tailored to the characteristics of benefit estimation tasks. In experiments involving sixteen models, our approach outperformed existing uncertainty quantification methods in the majority of cases. We also conducted index tuning tests on six datasets. By applying the Beauty framework, we eliminated worst-case scenarios and more than tripled the occurrence of best-case scenarios.
Index Terms:
Index tuning, database, benefit estimation, uncertainty quantification.
1 Introduction
Indexes are a crucial means for accelerating database queries. However, poorly configured indexes not only waste storage resources but can also degrade system performance. The automatic creation of suitable indexes has long been a major research focus, commonly referred to as index tuning [1, 2]. The component that addresses this challenge, known as the index advisor, typically consists of two parts. The first is the index configuration enumerator, which continuously generates potential index configurations. These configurations are then passed to the second component, the benefit estimator, responsible for estimating the benefits these index configurations would bring to the workload, without actually creating the indexes.
Benefit estimation (BE) plays a critical role in effective index tuning. The most common method involves using what-if tools [3] to create virtual indexes and utilizing the query optimizer to estimate their potential benefits. However, this approach has two significant drawbacks: (1) Low inference efficiency: The optimizer needs to sequentially compare the costs of almost all execution plans of a query. This process accounts for nearly 90 percent of the entire index tuning time [4]. (2) Low accuracy: The optimizer’s cardinality and cost estimators often produce significant errors, especially for complex queries.
To address these issues, researchers have studied learning-based models for benefit estimation, such as [5, 6, 7]. While these models generally improve efficiency, they present new challenges when deployed in real-world systems.
(1) Poor stability: Learning-based models typically outperform what-if tools on data similar to their training sets, but they can generate highly erroneous results when query patterns shift or training data is insufficient. As shown in Figure 1, we compare a learning-based BE method with a what-if tool on varying proportions of out-of-domain (OOD) samples. When the proportion reaches around 25%, the mean errors of two methods are comparable, but the mean error of the learning-based method becomes larger than the what-if tool as the OOD sample proportion increases.
Figure 1: Mean errors of a learning-based BE model and a what-if tool.
(2) Lack of interpretability: Learning-based models are often black-box systems, making them difficult to understand when and why they perform poorly in real-world applications.
(3) Complex management: These models require extensive data collection before deployment and ongoing monitoring to determine when updates are necessary, adding considerable overhead.
Due to the poor stability of learning-based models, some estimation results will inevitably be highly inaccurate and unreliable, leading the index advisor to select suboptimal indexes. Furthermore, the lack of interpretability prevents us from determining when or why the model produces unreliable results, making it difficult to fundamentally address the stability issue. In this paper, we propose a different approach by quantifying the uncertainty of the model’s results through specific metrics. When the model’s estimate for a test sample shows high uncertainty, the result is more likely to significantly deviate from the true value. This technique, known as uncertainty quantification (UQ), has been widely studied in fields where model errors have serious consequences, such as medical image detection [8], autonomous driving [9], and drug analysis [10].
By using this approach, we can return only the reliable benefit estimation results to the index advisor, while utilizing what-if tools to obtain the remaining results, thereby improving the reliability of the BE module. To achieve this, we propose the first uncertainty-aware BE framework, named Beauty (benefit estimation with uncertainty). It employs UQ techniques to combine the strengths of learning-based BE models and what-if tools, ensuring compatibility with existing systems. Additionally, by monitoring the proportion and causes of uncertain results during model runtime, we can determine when to update the model, reduce model management costs, and guide model design. We believe this framework will facilitate the broader adoption of learning-based models.
Although many UQ methods are designed for general tasks, their accuracy and efficiency are suboptimal when applied directly to BE tasks. To address this, we conducted a systematic study of the sources of uncertainty in learning-based BE models and performed a finer-grained analysis. Based on our findings, we propose a novel hybrid UQ method that integrates AutoEncoder and Monte Carlo Dropout into existing BE models. This method aligns with the structure of current learning-based estimators and outperforms general-purpose UQ methods in most scenarios. Additionally, we introduce a scheme that balances quantification accuracy and inference efficiency, tailored to the specific characteristics of the model. The source code and data for this study are available at  [^1] https://github.com/HIT-DB-Group/Beauty.
In summary, this paper makes the following contributions:
(1) We conducted a systematic study of the sources of uncertainty in learning-based BE models within the context of index tuning. Based on the characteristics of these tasks, we proposed a novel hybrid UQ method that allows for a flexible trade-off between efficiency and accuracy, depending on the specific characteristics of the BE task (Section 4). Our experiments demonstrate that the proposed UQ method outperforms general-purpose UQ methods in five out of six scenarios.
(2) We introduced Beauty, the first uncertainty-aware BE framework, which combines the strengths of learning-based models with what-if tools, enabling uncertainty-driven model updates. Our experiments across six benchmarks with varying budgets show that the Beauty framework significantly improves index tuning performance (Section 6).
(3) In our experiments, we applied existing UQ methods to various learning-based BE models, creating a total of 16 different models (Section 5). We thoroughly evaluated the impact of these methods on model accuracy and efficiency, confirming the superiority of our proposed UQ method.
2 Background
2.1 Index Advisors
Index advisors are responsible for index tuning, that is, recommending indexes for databases [11]. The typical process of index tuning is illustrated in Figure 2. An index advisor consists of three critical modules.
Figure 2: Typical structure of index advisors.
Workload Manager: The workload manager periodically collects a set of queries $\{q_{1},q_{2},\dots,q_{n}\}$ recently executed under the current index configuration $\mathcal{I}_{0}$ created on the database. The queries, along with their execution plans and weights (indicating frequencies or importance) $w_{i}\in\mathbb{R}$, form a workload $W$.
Index Configuration Enumerator: The index configuration enumerator generates potential index configurations $\mathcal{I}$ that aim to reduce the execution time of $W$ subject to the budget constraint such as the maximum storage space of $\mathcal{I}$. There are mainly two approaches to enumerating index configurations. The early research primarily focused on greedy algorithms [12, 13, 14] or linear programming [15]. The other approach models the problem as a reinforcement learning task, employing techniques such as Monte Carlo Tree Search [4], Multi-Armed Bandit [16, 17], or Deep Reinforcement Learning [18, 19].
Benefit Estimator: Let $c(q,\mathcal{I})$ denote the cost of executing query $q$ with the support of index configuration $\mathcal{I}$. The benefit of replacing $\mathcal{I}_{0}$ with $\mathcal{I}$ for query $q$ is defined as $B(q,\mathcal{I}_{0},\mathcal{I})=c(q,\mathcal{I}_{0})-c(q,\mathcal{I})$. For each enumerated configuration $\mathcal{I}$, the benefit estimator computes the estimated benefit of executing all queries in the workload $W$ without actually building the indexes in $\mathcal{I}$.
The index configuration enumerator and the benefit estimator operate interactively. Eventually, the index configuration $\mathcal{I}$ that maximizes the overall benefit
$B(W,\mathcal{I}_{0},\mathcal{I})=\sum_{i=1}^{n}w_{i}\cdot B(q_{i},\mathcal{I}_{0},\mathcal{I}).$ (1)
is recommended for the database. Clearly, maximizing $B(W,\mathcal{I}_{0},\mathcal{I})$ is equivalent to minimizing the total cost $c(W,\mathcal{I})=\sum_{i=1}^{n}w_{i}\cdot c(q_{i},\mathcal{I})$.
2.2 Learning-based Benefit Estimation (BE) Models
The benefit estimator plays a crucial role in index tuning as its efficiency directly impacts the overall performance of index tuning, and its accuracy affects the behavior of the index configuration enumerator [15, 20, 21]. Traditionally, what-if tools have been used to create hypothetical indexes, and the query optimizer is employed to estimate the total cost $c(W,\mathcal{I})$. However, due to their inefficiency and inaccuracy, particularly, for complex queries, machine learning (ML) techniques have been recently introduced to develop faster and more accurate BE models. One approach such as DISTILL [1] involves training multiple lightweight BE models for each query template to collectively estimate benefits. The other approach is to train a single complex BE model using historical query execution data, either to completely replace the what-if tool or to adjust the outputs of the what-if tool. The latter approach is the main focus of our study. LIB [7] and AMA [6] are notable examples of this approach. As shown in Figure 3, the typical architecture of a learning-based BE model consists of three modules.
Figure 3: The typical architecture of a learning-based BE model.
Feature Extractor: The feature extractor represents the information relevant to the BE model as a set of vectors for processing. The encoded information includes the index configuration $\mathcal{I}$, the SQL statement of query $q$, the query plan of $q$ under the current index configuration $\mathcal{I}_{0}$, the database schema, the database statistics, the hardware environment, and so on. According to the data type, each feature can be classified as numerical or categorical. The features are encoded into a set of vectors $V^{1}=\{\mathbf{v}_{1}^{1},\mathbf{v}_{2}^{1},\dots,\mathbf{v}_{t}^{1}\}$. Depending on the model design, the number $t$ of vectors in $V^{1}$ can be fixed (e.g., AMA [6]) or variable (e.g., LIB [7]) for different queries and environments. For each vector $\mathbf{v}_{i}^{1}$, the elements encoding the categorical features are replaced by higher-dimensional vectors using one-hot encoding or learnable word embedding, resulting in a vector $\mathbf{v}_{i}^{2}$ with higher dimensionality than $\mathbf{v}_{i}^{1}$. Thus, the input vector set for the BE model is $V^{2}=\{\mathbf{v}_{1}^{2},\mathbf{v}_{2}^{2},\dots,\mathbf{v}_{t}^{2}\}$.
Hidden Vector Encoder: Each vector in $V^{2}$ represents an aspect of the information relevant to the BE model. Since the vectors in $V^{2}$ are often sparse, they are usually consolidated into a compact and information-rich hidden vector $\mathbf{v}^{h}$. If the number of vectors in $V^{2}$ varies across queries, a recurrent neural network (RNN), a graph neural network (GNN), or the multi-head attention mechanism can be used as the encoder for $\mathbf{v}^{h}$, allowing each vector $\mathbf{v}_{i}^{2}\in V^{2}$ to interact with other vectors $\mathbf{v}_{j}^{2}\in V^{2}$. It first produces an intermediate representation $V^{h}=\text{Aggregate}(V^{2})=\{\mathbf{v}_{1}^{h},\mathbf{v}_{2}^{h},\dots,\mathbf{v}_{t}^{h}\}$, and then, the vectors in $V^{h}$ are processed into the final hidden vector $\mathbf{v}^{h}=\text{Pool}(V^{h})$ using techniques such as pooling or adding a super node. If the number of vectors in $V^{2}$ is fixed, a multi-layer perceptron (MLP) is often used as the encoder for the hidden vector, i.e., $\mathbf{v}^{h}=\text{MLP}(V^{2})$.
Predictor: The predictor, typically a MLP, takes the hidden vector $\mathbf{v}^{h}$ as input and generates the final output $\hat{y}=\text{MLP}(\mathbf{v}^{h})$ which represents an estimate of either $c(q,\mathcal{I})$ or $B(q,\mathcal{I}_{0},\mathcal{I})$.
2.3 Uncertainty in Machine Learning (ML) Models
Uncertainty is the quality or state of being uncertain, which arises in every stage of a ML task, including data acquisition, feature encoding, model design, model training, and model inference [22]. Naturally, uncertainty is inherent in ML-based index tuning, which will be explained in subsection 4.2. As illustrated in Figure 4, there are two different types of uncertainty, often referred to as aleatoric uncertainty and epistemic uncertainty [23].
Figure 4: A set of data points and a Bayesian neural network (BNN) trained based on them. The curves on the left represent the fitting curves for three different neural network samples from the BNN. The relationships between these curves and the data points demonstrate the distinction between aleatoric and epistemic uncertainty.
Aleatoric uncertainty, also known as data uncertainty or statistical uncertainty, refers to the notion of randomness, that is, the variability in the outcome of an experiment (i.e., random errors) caused by inherently random effects. For instance, the data collected by sensors usually has aleatoric uncertainty because the same input target might correspond to different sensing outputs at random. In principle, aleatoric uncertainty cannot be reduced by incorporating additional information such as collecting more data or using better models [24]. Aleatoric uncertainty can be further divided into homoscedastic and heteroscedastic depending on if the inherent randomness varies with the input target.
Epistemic uncertainty, also known as model uncertainty or systematic uncertainty, refers to uncertainty caused by the ignorance of knowledge (about the best model). Epistemic uncertainty may result in errors generated in the design, training and inference of a model. For example, if the inference of a model is made on a dataset that has a distribution significantly different from the model’s training dataset, the model might provide inaccurate outputs due to ignorance. As opposed to uncertainty due to randomness, uncertainty caused by ignorance can in principle be reduced based on additional information. In other words, epistemic uncertainty refers to the reducible part of the (total) uncertainty, whereas aleatoric uncertainty refers to the irreducible part.
Uncertainty Quantification (UQ): A number of model-independent and model-dependent methods have been proposed to quantify the two types of uncertainty. The first category includes Bayesian neural networks (BNNs) [25, 26], Monte Carlo dropout (MCD) methods [27], and ensemble methods [28, 29], which do not rely on the specific models in which the uncertainty is quantified. The second category incorporates UQ into model-specific structures. For example, HSTO [30] replaces the multi-head attention mechanism with hierarchical stochastic self-attention to achieve UQ. In this paper, we will reproduce these methods in the experimental evaluation and compare them with our proposed hybrid UQ approach.
3 Problem
Since collecting training data is costly, ML models inevitably encounter out-of-distribution (OOD) data, leading to unreliable outputs. The unreliable results produced by learning-based benefit estimators can lead index advisors to select inappropriate indexes, significantly degrading database performance. Therefore, as discussed in section 1, we will explore how UQ techniques can be leveraged to determine whether the results produced by learning-based BE models are reliable and how to integrate uncertainty-aware benefit estimators into index advisors. Specifically, this paper addresses the following three key research questions:
Q1: How can the uncertainty in a learning-based BE model be precisely quantified? Accurately quantifying the uncertainty in a learning-based BE model presents a significant challenge, and directly applying the existing methods often yields suboptimal performance. As shown in Figure 5, on a OOD dataset, model error and model uncertainty are generally not highly linearly correlated, that is, a large model error may not mean large model uncertainty. In this paper, the first problem is how to precisely quantify the uncertainty in a learning-based BE model. Rather than applying the existing general-purpose UQ methods directly, we propose a hybrid quantification method in section 4 tailored to the unique characteristics of the BE problem, providing more precise quantification of uncertainty than the general-purpose methods.
Figure 5: Joint distribution of model error and model uncertainty evaluated using UQ methods BNN [31], MCD [27], HSTO [30], and Ensemble [32]. $r$: Pearson’s correlation coefficient.
Q2: What impacts does UQ have on a BE model? In section 5, we will conduct an empirical study to understand the impacts of UQ on BE models by evaluating the accuracy, training cost, and inference cost of the BE models enhanced with UQ. In subsection 4.4, we will also explore a new method for determining when a BE model needs to be updated based on quantified uncertainty.
Q3: What impacts does a UQ-enhanced BE model have on the index advisor? In section 6, we will evaluate 8 UQ-enhanced BE models with 6 workloads across 36 scenarios to investigate the effects of the accuracy and efficiency of UQ on the index advisor.
Due to space constraints, we focus exclusively on BE models based on neural networks (NNs). NN-based models have been receiving increasing attention and generally achieve higher accuracy. The approach proposed in this paper can be easily extended to non-NN models by replacing the UQ methods for non-NN models such as [33, 34].
4 The Beauty Framework
To answer the questions in section 3, we propose our new framework for benefit estimation called Beauty (Benefit estimation with uncertainty).
4.1 Overview
The process of index tuning with Beauty is illustrated in Figure 6. After generating an index configuration by the enumerator, the advisor uses the uncertainty-aware learning-based BE model to estimate both the benefits that the configuration brings to the queries in the workload and the confidence levels of the estimated benefits. Then, the result filter examines the estimated benefits and their uncertainty with respect to a specified uncertainty threshold. The estimated benefits with uncertainty below the threshold are deemed reliable and are returned to the enumerator. For the remaining estimated benefits on which the BE model has low confidence due to insufficient training data or feature extraction deficiencies, the what-if tool that is usually more robust than learning-based BE models are invoked instead for benefit estimation.
Figure 6: The index tuning process with Beauty.
4.2 Uncertainty in Learning-based BE Models
The existing UQ methods are designed for general ML models and scenarios, lacking specific knowledge necessary for benefit estimation. To overcome this limitation, we propose a hybrid UQ method tailored to the unique characteristics of the BE problem, providing more precise quantification of uncertainty than the general-purpose methods. To this end, we conduct the first analysis of the sources of uncertainty in the BE problem and their relationships to the structures of learning-based BE models, which lays foundations for designing the uncertainty-aware BE model in Beauty.
Uncertainty emerges throughout the whole process of learning-based BE models from design to training and to inference. We divide the uncertainty into five categories which are summarized in Table I.
TABLE I: Five categories of uncertainty in learning-based BE models.
Category Type Cause Module for Quantification
U1 Epistemic Insufficient information modeling Predictor
U2 Epistemic Flawed feature extraction Predictor
U3 Epistemic Inadequate model design and training Hidden Vector Encoder
U4 Epistemic Insufficient training data Hidden Vector Encoder
U5 Aleatoric Randomness in query execution time Predictor
U1: Epistemic uncertainty due to insufficient information modeling. Estimating the benefit of an index configuration for a query is a complex task, as it requires invoking the query optimizer to retrieve the plans of the query before and after creating the indexes in the configuration. Existing learning-based estimators only model the information most relevant to this query such as the index configuration, the current query plan, and the plan chosen by the optimizer after virtually creating the indexes in the configuration. However, some overlooked factors can also significantly affect query execution costs. One such factor is data distributions which determine the cardinalities of intermediate results produced during the execution of a query. Existing learning-based estimators either do not model data distributions or model them in an over-simplified manner. Cardinality estimation itself is challenging and is a key research focus in the AI4DB field. State-of-the-art data-driven cardinality estimation methods are often built on deep-learning models which are more complex than those used in the existing BE models. Another factor that significantly affects query execution costs is the configuration of the database system such as the knobs work_mem and shared_buffers. Finally, the implementation of the query optimizer itself can alter the selected plan for a query and varies its execution cost. Currently, none of the learning-based estimators has taken all these relevant information into account. The ignorance of these information leads to epistemic uncertainty.
U2: Epistemic uncertainty caused by flawed feature extraction. Learning-based BE models require encoding various BE-related information into vectors, e.g., query statements, query plans, indexes, and database states. Incomplete encoding process may cause information loss. Moreover, flawed feature extraction techniques may represent different training samples into the same vector, resulting in a training set where the same sample attain conflicting labels, thus contributing to epistemic uncertainty.
U3: Epistemic uncertainty due to inadequate model design and training. A ML model’s expression power is limited by its structure and adopted mechanisms, as well as the settings of the hyperparameters and training process, e.g., batch size, optimization algorithm, learning rate, and stopping criteria. As a result, even on the training data, the model cannot be 100% accurate in general. The portion of the data that the model fails to fit yields epistemic uncertainty.
U4: Epistemic uncertainty caused by insufficient training data. Training a BE model requires collecting the execution plans and costs of a set of queries under different index configurations, which places a heavy burden on the system. Additionally, the collected data might not cover all application scenarios, and therefore, out-of-domain (OOD) issues often occur during inference. For instance, if a BE model was trained mostly with OLAP queries but is applied mostly to OLTP queries during inference, the model’s accuracy tends to decline in practice.
U5: Aleatoric uncertainty due to the randomness in query execution time. Due to physical factors such as caching, system resource scheduling, and hardware conditions, the execution time of a query is not fixed but a random variable within a certain range. Therefore, if a BE model estimates benefits based on observed query execution time, the model’s outputs must exhibit randomness, and the randomness is related to the characteristics of queries, resulting in heteroscedastic aleatoric uncertainty.
4.3 Hybrid Quantification Method
In this section, we propose a hybrid method for quantifying uncertainty in a learning-based BE model. We first describe which modules in the BE model is each category of uncertainty mainly quantified from, followed by the design of specific quantification methods.
4.3.1 Modules for Quantifying Uncertainty
As described in subsection 2.2, there are three main modules in a learning-based BE model: the feature extractor, the hidden vector encoder, and the predictor. Here, we analyze the relationship between the five categories of uncertainty listed in subsection 4.2 and the modules of the BE model, which lays the foundation for our hybrid UQ method.
Quantifying Uncertainty from Hidden Vector Encoder Module. Uncertainty U3 and U4 arise from limited model capacity and insufficient data, leading to the discrepancy between the benefit function fitted by the model and the true benefit function. Uncertainty U3 and U4 are unrelated to the feature extractor. The hidden vector encoder and the predictor are responsible for the mapping process, with the hidden vector encoder being the key module that determines the model’s fitting ability. It transforms rich, high-dimensional information into a low-dimensional space, and the quality of the hidden feature vector $\mathbf{v}^{h}$ it produces directly reflects the model’s ability to fit the benefit function. Consequently, uncertainty U3 and U4 is quantified in the hidden vector encoder module.
Quantifying Uncertainty from Predictor Module. Due to the presence of U5, the output of the BE model exhibits randomness. However, the current benefit estimator produces a fixed estimated benefit value given a fixed input $q$, $\mathcal{I}_{0}$ and $\mathcal{I}$, so we need to modify the BE model to make its output random, with a distribution as close as possible to the true benefit distribution. Among the two modules responsible for the mapping process, the hidden vector encoder primarily removes redundant information and extracts features, and it has already been used to quantify uncertainty U3 and U4. Therefore, the task of introducing randomness into the output is assigned to the predictor module which directly generates the output. Consequently, uncertainty U1, U2, and U5 is quantified in the predictor.
Uncertainties U1 and U2 arise from deficiencies in the feature extractor: some data with different benefits $B(q,\mathcal{I}_{0},\mathcal{I})$ result in identical input vectors $V^{1}$ for the model. For the model, this still implies that the output exhibits randomness, i.e., the same input corresponds to multiple possible outputs. Thus, uncertainty U1 and U2 can also be quantified within the predictor module.
4.3.2 Quantifying Uncertainty from Hidden Vector Encoder
Given a learning-based BE model, uncertainties U3 and U4 represent the model’s fitting error with respect to the benefit function $f$, which is reflected in the quality of the hidden feature vector $\mathbf{v}^{h}$ extracted by the hidden vector encoder module. To evaluate the quality of $\mathbf{v}^{h}$, we employ the well-known concept of AutoEncoder [35] and incorporate a decoder network structure into the BE model. The basic idea is as follows. The added decoder tries to reconstruct the input $V^{1}$ of the hidden vector encoder from the intermediate representation $V^{h}$, i.e., $\hat{V}^{1}=Decoder(V^{h})=\{\hat{\mathbf{v}}_{1}^{1},\hat{\mathbf{v}}_{2}^{1},\dots,\hat{\mathbf{v}}_{t}^{1}\}$. The quality of the reconstruction is assessed by the total mean squared error (MSE) between all pairs of corresponding vectors $\mathbf{v}_{i}\in V^{1}$ and $\hat{\mathbf{v}}_{i}\in\hat{V}^{1}$, that is,
$\mathcal{U}_{1}=\frac{1}{t}\sum_{i=1}^{t}MSE(\mathbf{v}_{i}^{1},\hat{\mathbf{v}}_{i}^{1}).$ (2)
$\mathcal{U}_{1}$ quantifies the uncertainty in the BE model with respect to the input $V^{1}$ of the hidden vector encoder module.
The rationale behind this approach lies in the fact that for the decoder to accurately reconstruct $V^{1}$ from $V^{h}$, two essential conditions must be satisfied. First, the structure of the hidden vector encoder must be appropriate. The network structure of the decoder should be symmetric to that of the encoder. Only if the encoder’s structure can gradually eliminate unimportant features and reduce dimensionality, the decoder can reverse this process and reconstruct $V^{1}$ based on the essential features. Second, the quality of the latent representation $V^{h}$ must be high. Given that the structure of the decoder is appropriate, the quality of the decoder’s input also determines the quality of the reconstruction. Therefore, when the divergence between $V^{1}$ and $\hat{V}^{1}$ is very small, the encoder has processed $V^{1}$ effectively, indicating low uncertainty in the hidden vector encoder module.
As depicted in Figure 7, the encoder transforms $V^{2}$ to $\mathbf{v}^{h}$, but the decoder transforms $V^{h}$ to $\hat{V}^{1}$. Obviously, the encoder and the decoder do not carry out strictly inverse processes. First, $V^{h}$ is used as the input of the decoder rather than $\mathbf{v}^{h}$. This is because the pooling in the encoder that transforms $V^{h}$ to $\mathbf{v}^{h}$ irreversibly compresses most information, e.g., the number of vectors in $V^{h}$. The loss of information makes subsequent reconstruction in the decoder impossible. Second, $V^{1}$ is set as the target of the reconstruction process carried out by the decoder instead of $V^{2}$. This is because after transforming the categorical features in the vectors in $V^{1}$ to embeddings, the sparsity of the vectors increases, and the richness of information decreases, which significantly increases the difficulty in reconstruction carried out by the decoder. Our experiments verify that the design of the decoder is reasonable and effective.
Figure 7: The Beauty framework for joint benefit estimation and uncertainty quantification.
In Eq. (2), the MSE between $\mathbf{v}^{1}_{i}$ and $\hat{\mathbf{v}}^{1}_{i}$ is computed, where all elements of $\mathbf{v}^{1}_{i}$ and $\hat{\mathbf{v}}^{1}_{i}$ are considered equally important. However, in practice, different features can have different impacts on BE, and important features can be assigned higher weights. In engineering practice, MSE can be replaced by weighted MSE, and the practitioners can properly set weights to the dimensions of $\mathbf{v}^{1}_{i}$ and $\hat{\mathbf{v}}^{1}_{i}$.
4.3.3 Quantifying Uncertainty from Predictor
After using the decoder to quantify uncertainty U3 and U4 in the hidden vector encoder, we proceed to quantify uncertainties U1, U2, and U5 in the predictor module. To ensure compatibility with the existing benefit estimator, we cannot assume that the model adopts specialized mechanisms, such as GNN or multi-head attention. Thus, considering the available UQ methods [22], we have two options:
(1) Bayesian Neural Networks (BNNs). BNNs have been extensively studied for UQ. Let $\mathcal{N}$ be a neural network (NN). To quantify the uncertainty in $\mathcal{N}$, a BNN $\mathcal{B}$ is induced, which has the same structure as $\mathcal{N}$ but treats each parameter $w$ in $\mathcal{N}$ as a random variable $X_{w}$ following a probability distribution. Let $W$ be the set of all parameters in $\mathcal{N}$ and $\mathbf{X}_{W}=\{X_{w}\mid w\in W\}$ be the set of corresponding random variables in $\mathcal{B}$. Given a training set $D=\{(\mathbf{x}_{i},y_{i})\mid 1\leq i\leq n\}$, the goal is to learn the posterior probability density function $p(\mathbf{X}_{W}|D)$, which is given by
$p(\mathbf{X}_{W}|D)$ $=\frac{p(D|\mathbf{X}_{W})p(\mathbf{X}_{W})}{p(D)}$
[…]
The prior distribution $p(\mathbf{X}_{W})$ should reflect prior beliefs about the distribution of the weights $W$ before observing any data. Solving this equation directly can be challenging, so variational inference (VI) is often employed [36, 26]. Blundell et al. [31] proposed the renowned Bayes by Backprop (BBB) algorithm, which is compatible with backpropagation and facilitates learning a probability distribution. We select the BBB algorithm as the baseline in the subsequent experimental evaluation.
The trained BNN $\mathcal{B}$ and its inference process are depicted in Figure 4. The uncertainty in the NN $\mathcal{N}$ with respect to a test record $\mathbf{x}$ is assessed as follows. First, a collection of parameter settings $\{W_{1},W_{2},\dots,W_{m}\}$ is obtained through independent Monte Carlo sampling from the posterior distribution $p(\mathbf{X}_{W}|D)$. By setting the parameters $W$ to each sampled configuration $W_{i}$, a specific instance $\mathcal{N}_{i}$ of $\mathcal{N}$ is obtained. Given the test record $\mathbf{x}$ as input, let $y_{i}$ be the output of $\mathcal{N}_{i}$. The mean of $\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{m}$ is the prediction for $\mathbf{x}$, and the variance of $\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{m}$ is the uncertainty of $\mathcal{N}$ with respect to $\mathbf{x}$. In practice, the sampling and inference processes can be combined and performed in an interleaved, layer-by-layer manner on $\mathcal{N}$.
(2) Monte Carlo Dropout (MCD) Methods. Dropout layers are commonly used in deep NNs to prevent overfitting. During training, a dropout layer randomly sets elements of the input tensor to 0 with probability $p\in[0,1]$ and scales up the remaining elements by $1/(1-p)$ to maintain the sum over all inputs. The zeroed elements are chosen independently for each forward pass. During inference, dropout layers are typically disabled. However, to represent model uncertainty, dropout layers can be activated during inference. Gal and Ghahramani [27] proved that dropouts can approximate Bayesian inference in BNNs. Given a test record $\mathbf{x}$, model inference is repeated multiple times independently, yielding outputs $\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{m}$. The mean of these outputs is the prediction for $\mathbf{x}$, and the variance represents the model’s uncertainty with respect to $\mathbf{x}$.
In the BE scenario, we prefer the MCD method over BNNs for two reasons. First, BNNs require more extensive modifications to the BE model compared to MCD. Second, the training cost of a BNN is usually higher because BNNs often require multiple parameters to represent the distribution of a single neuron parameter. For instance, if a neuron’s weight follows a normal distribution $N(\mu,\sigma)$, two parameters, $\mu$ and $\sigma$, are needed to characterize the distribution. Thus, a BNN requires significantly more parameters than a standard NN, leading to increased training costs.
Specifically, we introduce MCD into the original MLP of the Predictor to approximate BNNs. After the encoder produces the hidden representation $\mathbf{v}^{h}$, a set of predictions $\{\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{m}\}$ is generated by the approximated BNNs, i.e., $\{\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{m}\}=MLP_{MCD}(\mathbf{v}^{h})$. The mean of $\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{m}$ is returned as the prediction for $B(q,\mathcal{I}_{0},\mathcal{I})$, and the uncertainty is quantified as
$\mathcal{U}_{2}=\textrm{Var}(\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{m}).$ (3)
4.3.4 Result Filter
Once the BE model with UQ capability (referred to as the BE-UQ model) provides the estimated benefit $\hat{y}$ and the quantified uncertainties $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, the Result Filter module (as shown in Figure 6) determines whether $\hat{y}$ should be returned to the Estimator.
Thresholds $\theta_{1}$ and $\theta_{2}$ are predefined for $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, serving as reliability indicators. If $\mathcal{U}_{1}$ exceeds the threshold $\theta_{1}$—associated with uncertainties U3 and U4—this suggests that $\hat{B}(q,\mathcal{I}_{0},\mathcal{I})$ likely diverges significantly from true value. Similarly, if $\mathcal{U}_{2}$ exceeds $\theta_{2}$—associated with uncertainties U1, U2, and U5—this indicates that $\hat{y}$ has a high variance. Specifically, the returned result $\hat{B}(q,\mathcal{I}_{0},\mathcal{I})$ should follow:
$\hat{B}(q,\mathcal{I}_{0},\mathcal{I})=\begin{cases}\hat{y}&\text{if }\mathcal{U}_{1}\leq\theta_{1}\ \text{and}\ \mathcal{U}_{2}\leq\theta_{2},\\ \text{What-if tools}&\text{otherwise}.\end{cases}$
The domain of $\theta_{1}$ and $\theta_{2}$ is $[0,+\infty)$, and their values vary based on factors such as model structure, hyperparameters, query workload, and the database instance. Setting fixed, universal thresholds for $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ is impractical. Therefore, we adopt a context-dependent approach to define relative thresholds $\theta_{1}$ and $\theta_{2}$: for a new sample, if the model’s uncertainty regarding it is comparable to the uncertainty observed during training when the model performed optimally with minimal uncertainty, the result is deemed reliable. Specifically, after training the BE-UQ model, we perform inference for each training sample $(q,\mathcal{I}_{0},\mathcal{I})$ and gather the quantified uncertainties $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$. The 90th percentile of $\mathcal{U}_{1}$ (or $\mathcal{U}_{2}$) across all training samples is designated as $\theta_{1}$ (or $\theta_{2}$).
4.4 Model Deployment
4.4.1 Model Training
Compared to the original BE model, our BE-UQ model incorporates two additional modules: the decoder and MCD. Since the decoder cannot be trained concurrently with the original BE model, a redesigned training process is required. In our index advisor, the BE-UQ model is trained in two distinct phases. First, the BE model—which includes the feature extractor, hidden vector encoder, and Predictor—is trained following the original BE model’s method. In the second phase, the parameters of the feature extractor and hidden vector encoder are frozen, and the decoder is trained by minimizing the MSE loss.
The rationale behind this training method is that BE remains the primary objective of the BE-UQ model, while UQ is secondary. Our approach ensures that the accuracy of the BE model is maintained, regardless of whether UQ is utilized. By preserving the accuracy of the hidden vector encoder, we ensure that the input information is effectively encoded in the intermediate representation $V^{h}$. This, in turn, enables the decoder to accurately reconstruct the input vector set $\hat{V}^{1}$ and correctly quantify uncertainty.
Although there are alternative ways to train the BE-UQ model, such as jointly training the hidden vector encoder and decoder followed by training the Predictor, this method shifts the encoder’s focus toward UQ rather than BE, which conflicts with our objectives. Therefore, our proposed training method is better suited for the index tuning task.
4.4.2 Uncertainty-Driven Model Updating
The performance of the BE-UQ model may decline when the database or query workload changes. To maintain the accuracy and stability of the model, it must be updated promptly when performance starts to decrease. There are two common ways to determine whether a machine learning model requires updating.
The first category samples predictions during model inference and compares them with actual values. A model update is needed if the error is significantly large. However, this approach has two main drawbacks. First, the actual benefits $B(q,\mathcal{I}_{0},\mathcal{I})$ for the sampled instances $(q,\mathcal{I}_{0},\mathcal{I})$ must be known. To determine $B(q,\mathcal{I}_{0},\mathcal{I})$, the indexes in $\mathcal{I}$ must be created, and the query $q$ must be executed based on those indexes, which incurs substantial time and space costs. Second, this method may cause delayed updates due to sampling errors.
The second category triggers model updates when a decrease in system efficiency is detected. However, this approach causes updates to lag behind the onset of performance regression. More importantly, in the context of index tuning, the optimal index is unknown. Even if an erroneous BE result leads to the creation of a suboptimal index, system performance may still improve slightly. Thus, it is difficult to determine whether the model requires updating based on this result.
In the Beauty framework, we have a more accurate and timely metric to determine whether an update is needed: if the what-if tool estimates benefits for a significant portion of test samples $(q,\mathcal{I}_{0},\mathcal{I})$, this likely indicates that the BE-UQ model no longer fits the current dataset, workload, or system conditions and requires updating. Additionally, the uncertainties $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ returned by the model can help guide the update process.
(1) If the what-if tool estimates the benefits for a significant portion of test samples $(q,\mathcal{I}_{0},\mathcal{I})$ due to $\mathcal{U}_{1}>\theta_{1}$, this suggests that the BE-UQ model cannot provide effective latent vectors under the current workload. For example, if there have been substantial changes in query patterns or index configurations, the BE-UQ model may fail to produce accurate results. In this case, more data must be collected from the current workload to retrain or fine-tune the model.
(2) If the what-if tool estimates the benefits for a significant portion of test instances $(q,\mathcal{I}_{0},\mathcal{I})$ due to $\mathcal{U}_{2}>\theta_{2}$, this indicates that the BE-UQ model is unable to distinguish between too many queries and index configurations in the current workload. Due to insufficient modeling information or flaws in feature extraction, the model estimates a wide range of benefits for the same test instance, leading to unreliable predictions. In such cases, the designer must identify the missing information or flaws in feature extraction and redesign the model based on the current workload.
4.4.3 Improving Efficiency.
During index tuning, the efficiency of benefit estimation significantly affects the tuning time and, in some cases, can influence the final tuning quality, particularly with algorithms like DTA that can be interrupted at any point. Faster estimation allows for exploring more index configurations, ultimately leading to better results. Although our framework enhances the reliability of benefit estimation, it inevitably introduces additional overhead. Here, we present a more efficient version for scenarios with tight time constraints. In terms of quantification overhead, $\mathcal{U}_{2}$ incurs greater cost than $\mathcal{U}_{1}$: the AutoEncoder requires an extra decoder inference, while MCD necessitates dozens of inferences by the Prediction Module to ensure result stability [37].
In terms of the importance of quantification, $\mathcal{U}_{1}$ is more critical than $\mathcal{U}_{2}$: (1) Among the uncertainties affecting $\mathcal{U}_{2}$, U1 and U2 arise from flaws in the model design. These can be mitigated by increasing the diversity of encoded information and designing a more effective feature extraction module. U5 is an aleatoric uncertainty that cannot be eliminated due to the inherent randomness of query execution. However, in practice, if the same query is executed multiple times, the average execution time can be used as the estimation target. (2) The uncertainties affecting $\mathcal{U}_{1}$, namely U3 and U4, cannot be completely resolved. It is difficult for the system to gather all possible combinations of queries and index configurations, and the model cannot achieve perfect accuracy on the training data. Consequently, the trained model will inevitably encounter workloads that deviate from those seen during training.
Therefore, when the feature extraction module is effective and the encoded information is sufficient, $\mathcal{U}_{2}$ quantification can be disabled. While this may slightly reduce the accuracy of the UQ, it significantly improves the inference efficiency of the BE-UQ model.
5 Empirical Study on Uncertainty Quantification
In this section, we compare different UQ methods from three aspects, namely the UQ capability, the efficiency and the impact on the accuracy of BE.
5.1 Datasets Construction
Three famous database benchmarks are adopted in the empirical study. Table II summarizes the databases and the workloads.
TPC-H: Following Leis et al. [38, 39], we randomly generated 30 queries from each template in TPC-H, excluding templates 2, 17, and 20. These were excluded because their queries have significantly longer execution times, which would disproportionately affect the overall workload cost. Indexes accelerating these queries are typically preferred over those for other templates.
TPC-DS: We also randomly generated 30 queries from each template in TPC-DS. Templates 4, 6, 9, 10, 11, 32, 35, 41, and 95 were excluded for the same reason given above.
JOB: This workload consists of 30 queries randomly generated from each template in JOB.
To increase the complexity of index tuning, we ingested synthetic queries into these workloads using the method devised by Yu et al. [39]. The enhanced workloads are referred to as TPC-H+, TPC-DS+ and JOB+, respectively.
TABLE II: Summary of databases and workloads and the maximum runtime for index tuning algorithm DTA.
Benchmark DB SF/Size # Queries # Templates # Synthetic Queries Avg. # Plan Nodes Maximum Runtime (mins)
TPC-H SF = 10 570 19 0 14 20
TPC-DS SF = 1 2700 90 0 29 45
JOB 13GB 3390 113 0 24 60
TPC-H+ SF = 10 950 19 380 10 30
TPC-DS+ SF = 1 4500 90 1800 19 60
JOB+ 13GB 5650 113 2260 16 75
To the best of our knowledge, no systematic experiments have previously explored uncertainty involved in index tuning tasks. To investigate the impacts of UQ on the accuracy and the efficiency of learning-based benefit estimators, as well as their ability to quantify uncertainty, we designed the following experimental methodology:
We ran the Autoadmin and Extend algorithms on three distinct workloads: TPC-H+, TPC-DS+, and JOB+, under various budget constraints to collect index configurations for tuning. After materializing these configurations, we gathered query execution data, represented as [(query, index config) $\rightarrow$ benefit], and formed a comprehensive dataset. This dataset was then divided into three segments based on query templates and indexes:
Training Dataset: Used for model training, containing approximately 50% of the total data. This subset only includes some of the query templates and indexes.
Test Dataset:Comprises about 15% of the total data, where each query or index configuration appeared in the training set but was not simultaneously seen by the model. This segment tests the model’s generalization capability within in-domain data.
Uncertainty Dataset:This dataset accounts for about 35% of the total data. It includes queries and indexes that are not present in the first two datasets, and no similar queries (from the same template) appear. For each model mentioned in Subsubsection 5.2.2, we will test using these samples, excluding those with errors below the 90th percentile of the Training Dataset errors. This way, the remaining samples are those the model fails to predict correctly due to U1-U4, and the model should assign a high uncertainty value to them.These samples are designated as Uncertainty Samples.
There are two points to note. First, the samples that different models fail to predict accurately vary, resulting in an inconsistent number of Uncertainty Samples, but the difference does not exceed 10%. Second, we use query costs to calculate the benefit rather than the execution time of the query, so we do not include U5-type uncertainty samples here.
5.2 Experimental Setup
5.2.1 BE Models
In this experiment, we adopt two well-known learning-based BE models:
(1) LIB. The BE model proposed in [7] encodes the operators and information in a query plan that can be influenced by indexes, forming a variable-length Index Optimizable Operations Set. After encoding through Multi-head Attention, the pooling operation retrieves $v^{h}$, culminating in the output of the BE via a MLP.
(2) AMA. The model in [6] utilizes multiple channels to extract information and employs a weighted sum to compare query plans, selecting the faster option. Following [7], we replaced AMA’s classification layer with an MLP for benefit estimation.
It is important to note that while we aim to maintain consistency with the original paper regarding the two BE models, complete consistency may not be achievable. Our primary focus is on studying the impact of the UQ method rather than directly comparing the two BE models.
5.2.2 UQ Methods
We integrate five types of UQ methods into the aforementioned BE models, forming sixteen distinct BE-UQ models for our experiment:
(1) Ours. This model integrates the hybrid quantification method proposed in this paper, resulting in LIB-Hybrid and AMA-Hybrid. We also compare an alternative version that utilizes efficiency enhancements, represented as LIB-AE and AMA-AE.
(2) BNN. We replace the MLP in the Prediction Module of both models with BNN, forming LIB-BNN and AMA-BNN.
(3) MCD. We compare three variants of this method in the experiment [40]: the first variant replaces all dropout layers in the model with MCD, resulting in LIB-MCD-all and AMA-MCD-all. The second variant applies MCD only to the last layer of the model’s Prediction Module, forming LIB-MCD-last and AMA-MCD-last. The third variant adopts the approach proposed in [37], which utilizes determinantal point processes for sampling to enhance diversity with minimal inference time increase, forming LIB-MCD-dpp and AMA-MCD-dpp.
In addition to the above methods, we also compare two other types of UQ methods.
(4) Ensemble. Ensemble methods refer to machine learning techniques such as Bagging and Boosting, which create multiple models and then combine them to produce improved predictions. Ensemble methods generally yield more accurate results than any single model. Lakshminarayanan et al. [32] utilize ensemble methods to quantify model uncertainty. A set of models can be induced either independently (e.g., in Bagging) or dependently (e.g., in Boosting). Given a test record $\mathbf{x}$ as input, these models produce outputs $\hat{y_{1}},\hat{y_{2}},\dots,\hat{y_{m}}$, respectively. The mean of $\hat{y_{1}},\hat{y_{2}},\dots,\hat{y_{m}}$ serves as the prediction for $\mathbf{x}$, while the variance of $\hat{y_{1}},\hat{y_{2}},\dots,\hat{y_{m}}$ indicates the uncertainty of the ensemble model regarding $\mathbf{x}$.
During hyperparameter selection, we retain the five best performing models for LIB and AMA, creating the LIB-Ensemble and AMA-Ensemble.
(5) HSTO. [30] proposed Hierarchical Stochastic Self-Attention (HSTO) to replace the existing Self-Attention mechanism, enabling all models using Self-Attention to quantify uncertainty. This method utilizes the Gumbel-Softmax to transform the deterministic attention distribution for values in each attention head into a stochastic one. The prediction and uncertainty for a given test record $\mathbf{x}$ are obtained similarly to the MCD method.
Since this method relies on the Self-Attention module within the model, it is applicable only to LIB. We adopt the two versions mentioned in the original paper, resulting in LIB-STO and LIB-HSTO.
5.2.3 Implementation
We set the relative benefit $rb=\frac{B(q,\mathcal{I}_{0},\mathcal{I})}{c(q,\mathcal{I}_{0})}$ as the prediction target. To maintain compatibility with the LIB framework, we exclude data entries that exhibit negative benefits from our dataset. For each model discussed, we generate thirty sets of hyperparameters and train on the training dataset. Subsequently, we retain the model that demonstrates the best performance based on evaluations using the test dataset.
When quantifying uncertainty, we use the variance of results from five models as the quantification of uncertainty for the ensemble-based model. For other models, we perform inference 20 times [37] for each batch sample, using variance as the quantification of uncertainty.
(a) TPC-H+
(b) TPC-DS+
(c) JOB+
Figure 8: Uncertainty distribution of models using BNN, MCD, Ensemble, and HSTO.
Figure 9: Uncertainty distribution of models using Hybrid.
Figure 10: Uncertainty distribution of models using AE.
5.2.4 Hardware & Software
The experiments were carried out on a Ubuntu server with two Intel Xeon 4210R CPUs (10 cores, 2.40GHz), 256GB of main memory, and one NVIDIA GeForce RTX 3060 GPU. The DBMS is PostgreSQL 12.13.
5.3 Results on Uncertainty Quantification
We present Raincloud plots illustrating the uncertainty distributions for sixteen models across various benchmarks. Figure 8 displays the plots for models based on BNN, MCD, Ensemble, and HSTO. Figures 9 and 10 show the plots for models based on Hybrid and AE, respectively. Each subplot depicts the uncertainty distribution of samples as calculated by the model for the training dataset, test dataset, and Uncertainty Samples. Consequently, the greater the uncertainty of the orange samples representing Uncertainty Samples in the graph and the higher their differentiation from other samples, the more precisely the model’s uncertainty modeling is demonstrated, enabling it to effectively distinguish uncertainty data. The red lines in the graphs represent the average uncertainties across the three datasets.
TABLE III: Percentage of Uncertainty Samples exceeding the uncertainty threshold set by the Training Dataset.
LIB-based Model TPC-H+ TPC-DS+ JOB+ AMA-based Model TPC-H+ TPC-DS+ JOB+
BNN 13.03 40.20 16.04 BNN 45.26 65.32 53.67
MCD-all 68.08 40.98 26.22 MCD-all 51.73 54.17 45.61
MCD-last 61.79 29.32 5.92 MCD-last 21.76 21.27 22.11
MCD-dpp 16.02 36.91 8.35 MCD-dpp 78.28 34.83 30.76
Ensemble 79.01 53.44 29.32 Ensemble 38.73 77.62 60.22
AE 71.85 31.13 30.93 AE 92.68 84.19 61.00
Hybrid 81.80 57.65 32.88 Hybrid 91.74 78.13 47.82
$\mathcal{U}_{1}$ 51.88 27.21 20.68 $\mathcal{U}_{1}$ 88.87 54.21 20.00
$\mathcal{U}_{2}$ 59.70 43.46 20.76 $\mathcal{U}_{2}$ 54.93 47.50 34.87
STO 68.41 34.86 23.51
HSTO 63.64 35.13 24.58
Table III further illustrates the filtering capabilities of various methods for Uncertainty Samples. Since Hybrid includes two UQ modules, we adopt the following approach to ensure comparability with other models: for $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, we select the 90th percentile of the uncertainty distribution from the Training Dataset as their corresponding thresholds. We calculate the proportion $P$ of samples in the Training Dataset with uncertainty values lower than these thresholds. If $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ are probabilistically independent, this proportion should be 81%. For other models with only one UQ module, we employ the $P$-th percentile of the uncertainty distribution from the Training Dataset as a threshold. The table shows the proportion of samples in the Uncertainty Samples that exceed this threshold. The higher the proportion, the stronger the model’s ability to identify uncertain samples. For each benchmark in the table, we highlight the model with the most effective filtering in green and the second most effective in blue. Note that $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, as parts of the Hybrid, are excluded from this ranking.
From these results, we can draw the following conclusions:
First, the AE and Hybrid models consistently perform the best, followed by the Ensemble method. The other methods lag significantly behind in both stability and optimal performance in UQ. The best performance is seen with the AMA-AE model under the TPC-H+, which can identify 92.68% of Uncertainty Samples while misclassifying about 19% of training samples, significantly outperforming general-purpose methods. In more than two-thirds of the benchmarks, the AE and Hybrid models consistently rank among the top two, demonstrating better stability than the Ensemble method.
Second, the quality of uncertainty quantification is related to the base model and dataset. Across three benchmarks, the uncertainty modeling quality of the AMA base model consistently surpasses that of the LIB model, particularly evident in the proposed AE and Hybrid models. This may be attributed to the fact that the model’s ability to fit the data also determines its accuracy and capability in UQ. As illustrated in Table V, AMA’s accuracy is significantly higher than LIB’s. Accurate results necessitate effective encoding of inputs into a hidden vector, the foundational principle of our methodology, which accounts for these observations.
Overall, these 16 models demonstrate superior performance on the TPC-H+ bencamark compared to TPC-DS+, with the poorest performance observed on the JOB+ benchmark, where the most effective UQ methods filter only 61% of Uncertainty Samples.
Third, the settings for uncertainty thresholds significantly impact the ability to detect Uncertainty Samples. Uncertainty distribution within a dataset is usually confined to a narrow range, and the relationship between the threshold and this range dictates the model’s detection capabilities. For example, in the AMA-Ensemble model under TPC-H+, setting a filtering threshold at the 81th percentile of the training dataset excludes only 38.73% of Uncertainty Samples, whereas a threshold at the 75th percentile increases exclusion to 73.34%.
Fourth, HSTO, specifically designed for Self-Attention, did not demonstrate improved performance on this task. Neither model achieved optimal performance across the three benchmarks.
Fifth, among the three MCD-based methods, MCD-all outperforms the others, achieving optimal results in five out of six scenarios. MCD-dpp provides a richer sampling diversity compared to MCD-last and outperforms it in the majority of scenarios.
TABLE IV: Time for model inference per batch (ms).
Model TPC-H+ TPC-DS+ JOB+ Model TPC-H+ TPC-DS+ JOB+
LIB 2.76 2.30 5.58 AMA 0.18 0.18 0.25
BNN 61.85 44.46 50.75 BNN 10.77 7.41 10.54
MCD-all 184.59 45.03 155.14 MCD-all 6.19 2.94 4.90
MCD-last 60.00 89.63 133.04 MCD-last 5.23 4.13 5.23
MCD-dpp 60.01 147.86 83.63 MCD-dpp 260.82 78.39 51.45
Ensemble 22.72 14.15 29.90 Ensemble 1.82 0.80 1.71
AE 3.08 4.76 6.37 AE 0.52 0.51 0.51
Hybrid 60.74 103.95 98.16 Hybrid 10.28 9.82 9.26
STO 151.40 137.33 157.01
HSTO 64.46 109.25 327.03
TABLE V: The impact of uncertainty quantification methods on the prediction errors. Models that have the most negative impact are marked in red, while those with the most positive impact are marked in green.
Dataset LIB BNN MCD-all MCD-last MCD-dpp Ensemble STO HSTO AE Hybrid AMA BNN MCD-all MCD-last MCD-dpp Ensemble AE Hybrid
TPC-H+ 50th 0.13 +0.10 +0.01 +0.02 +0.04 +0.00 +0.03 +0.00 +0.09 +0.05 0.01 +0.02 +0.00 +0.00 +0.02 +0.02 +0.00 +0.00
95th 0.47 +0.02 +0.00 +0.00 -0.04 +0.00 +0.00 +0.00 +0.05 +0.01 0.05 +0.06 +0.01 +0.03 +0.00 +0.03 +0.00 +0.02
99th 0.65 -0.11 -0.05 -0.04 -0.07 -0.01 +0.00 -0.05 -0.12 -0.01 0.11 +0.07 +0.00 +0.00 +0.03 +0.01 -0.03 +0.02
TPC-DS+ 50th 0.06 +0.01 +0.00 -0.01 +0.00 +0.00 +0.00 +0.00 +0.00 +0.00 0.01 +0.01 +0.00 +0.00 +0.00 +0.00 +0.00 +0.00
95th 0.44 +0.03 +0.00 -0.02 +0.01 +0.00 +0.00 -0.02 +0.00 -0.04 0.20 +0.06 +0.02 +0.00 +0.09 +0.03 +0.01 +0.01
99th 0.71 +0.00 +0.00 +0.03 +0.00 +0.00 +0.00 +0.00 +0.02 +0.00 0.34 +0.08 +0.06 +0.05 +0.07 +0.04 +0.05 +0.07
JOB+ 50th 0.19 +0.13 +0.01 +0.03 +0.04 +0.00 +0.04 +0.05 +0.03 +0.01 0.03 +0.02 +0.00 +0.00 +0.05 +0.00 +0.00 +0.01
95th 0.57 +0.04 +0.00 +0.00 +0.00 +0.00 +0.02 +0.00 +0.03 +0.00 0.19 +0.06 -0.01 +0.00 +0.05 +0.00 +0.00 +0.04
99th 0.76 +0.16 -0.02 +0.02 +0.00 -0.02 +0.07 +0.09 +0.04 +0.04 0.37 +0.17 +0.00 +0.00 +0.11 +0.00 +0.08 +0.10
5.4 Results on Prediction Accuracy and Inference Efficiency
Table IV shows the time taken by different models to process a single batch, which includes predicting the relative error $\hat{rb}_{predicted}$ and estimating uncertainty, in milliseconds. Models in which the inference time with UQ is on the same order of magnitude as the base model are marked in green.
The primary factors influencing inference efficiency are the base model structure, model hyperparameters, and methods for UQ. LIB, utilizing heavier Multi-head Attention, requires significantly more inference time, up to an order of magnitude higher than AMA.For the models based on the same BE model, the optimal hyperparameters vary across different datasets, which can also lead to several times difference in inference efficiency. Methods like BNN, MCD, and HSTO necessitate multiple inferences on the same model, resulting in inference times that are proportionally related to the number of inferences, significantly exceeding the base model and AE by one to two orders of magnitude. Specifically, MCD-dpp introduces extra computational overhead due to sampling processes, notably affecting the faster base model AMA, and to a lesser extent, the heavier LIB model where it is less of a determining efficiency factor. The inference time for Ensemble models scales with the number of models involved; selecting five models during hyperparameter tuning has resulted in higher overall efficiency compared to other multiple-inference UQ methods. Of all methods evaluated, AE proves the most efficient, necessitating only an additional inference for the decoder network, which keeps its inference times close to those of the base model on all six benchmarks.
Table V evaluates the impact of incorporating UQ methods into BE models. We assess performance using the absolute difference between predicted and actual values, i.e. $|rb_{actual}-\hat{rb}_{predicted}|$. The table shows the 50th, 95th, and 99th percentiles of error distributions for LIB and AMA. For models to which UQ methods have been added, the table shows the change in error relative to the base model, where models that have the most negative impact are marked in red, and those with the most positive impact are marked in green.
Among all UQ methods, only BNN significantly affected the accuracy of the BE model, while the other methods had no substantial impact. Each model was retrained using 30 distinct sets of hyperparameters, so any minor variations in accuracy can be attributed to differences in hyperparameter settings. In the case of BNN, however, each neuron is replaced by a value sampled from a distribution, which effectively doubles the number of network parameters (as each deterministic neuron corresponds to values such as mean, variance, etc., in BNN) and significantly expands the hyperparameter space (e.g., requiring a prior distribution). This makes the training process fundamentally different and more complex, requiring more intensive hyperparameter tuning.
Comparing to the MCD-last, MCD-dpp had a higher effect on accuracy when AMA was used as the base model, more inference burden while more UQ capability, making it better suited for integration with heavier base models. , which might be related to the changes it made in the sampling process during inference. However, more research is needed to provide a definitive conclusion, which is beyond the scope of this study.
Among the other methods, since the AE method did not alter the original structure and parameters of the BE model after training, it theoretically should not affect the accuracy of the BE model.
6 Empirical Study on Index Tuning
6.1 Experimental Setup
In this section, we explore whether modeling uncertainty can enhance index tuning results. Therefore, we integrated the aforementioned BE model into the Beauty framework and performed index tuning on six benchmarks under different budget constraints.
6.1.1 Implementation
Leis et al. [38] show that no single index configuration enumeration algorithm delivers optimal performance in all scenarios. In our experiments, we used DTA [14], which balances runtime and solution quality well across various scenarios. This algorithm has a set maximum runtime and stops after enumerating all configurations or reaching this limit. However, following Leis et al.’s [38] implementation, the algorithm completes the current seed’s configuration even if it exceeds the time limit, occasionally surpassing the maximum runtime. We adjust the maximum runtime based on workload size, as shown in Table II.
Based on the performance of various models mentioned in Section 5, we selected nine models in our experiment: LIB, LIB-Ensemble, LIB-AE, LIB-Hybrid, AMA, AMA-Ensemble, AMA-AE, AMA-Hybrid, and what-if. The thresholds for each model were set as described in Subsection 5.3. When the uncertainty value of the benefits estimated by the BE-UQ model for a sample exceeds the threshold, we use what-if tools to estimate the benefits of that sample.
6.1.2 Evaluation Metrics
We use specific metrics to assess the performance of index tuning. Given a workload $W$, $\mathcal{I}_{0}$ represents the initial set of indexes, and $\mathcal{I}$ the set of selected indexes. We measure the quality of replacing $\mathcal{I}_{0}$ with $\mathcal{I}$ by the relative improvement in $W$’s cost: $\max\left(0,\frac{B(W,\mathcal{I}_{0},\mathcal{I})}{c(W,\mathcal{I}_{0})}\right)=\max\left(0,1-\frac{c(W,\mathcal{I})}{c(W,\mathcal{I}_{0})}\right).$
On some occasions, if the selected indexes increase the cost, they must be discarded, and thus the improvement in cost is zero.
Figure 11: Cost improvements achieved by different benefit estimation models.
TABLE VI: Rankings of cost improvements achieved by models.
Method # 1st Place # 2nd Place # 3rd Place # Worst
What-if 8 8 11 1
LIB 0 0 1 10
LIB-Ensemble 1 4 4 3
LIB-AE 3 1 1 4
LIB-Hybrid 3 7 0 4
AMA 3 3 3 6
AMA-Ensemble 4 2 6 7
AMA-AE 10 7 5 0
AMA-Hybrid 4 4 5 0
6.2 Results on Index Tuning
Figure 11 illustrates the index tuning results of nine methods under various budgets across six benchmarks. Table VI presents detailed statistics of cost improvements, highlighting the best performances in green and the second best in blue. From these results, we can draw the following conclusions:
First, integrating the UQ module into the two base models significantly improved index tuning results. The LIB model has seen a significant reduction in the number of worst-case scenarios, while the AMA model has experienced a substantial increase in the number of optimal configurations. Additionally, when integrating AE and Hybrid into the AMA model, the number of worst-case scenarios decreased from six to zero.
Second, the performance of the integrated Hybrid model is superior to that of the Ensemble-based model. As shown in Table III, Hybrid’s ability to quantify uncertainty surpasses that of Ensemble, allowing it to more accurately identify the parts of the learning-based BE model that it cannot accurately estimate, thereby improving BE’s accuracy. The DTA algorithm requires a specified maximum runtime, and usually, it cannot enumerate all index configurations by the end. Therefore, the greater the delay in BE-UQ inference, the fewer configurations the index advisor can explore. As seen in Table IV, Hybrid has a higher delay. Thus, by providing more accurate uncertainty quantifications, Hybrid finds better results with fewer configurations.
Third, AE’s results are slightly better than Hybrid’s. The accuracy of the LIB base model is not high, and in this situation, Hybrid’s UQ accuracy is higher than AE’s, enabling it to identify as many uncertain data points as possible, resulting in better performance than AE. Conversely, when using AMA as the base model, AE achieves higher identification accuracy and lower inference delay, allowing for more accurate enumeration of more index configurations. Therefore, it achieves the highest number of optimal configurations and surpasses the what-if tools. For index tuning tasks, Hybrid is preferable when there is sufficient time available for tuning. If time is limited and the BE model demonstrates high accuracy, selecting AE can result in a better index configuration.
Fourth, in scenarios abundant with Uncertainty Samples, the what-if approach remains a stable and precise method, achieving the best or nearly the best results across all six benchmarks. This consistency forms the foundation of our study, which aims to model the uncertainty in learning-based models, combining their accuracy and efficiency with the reliability of what-if tools.
7 Related Work
We presented the related work on UQ in subsection 2.3 and subsection 5.2. The related work on learning-based BE models is detailed in subsection 2.2. In this section, we introduce additional related work on index tuning.
Index Configuration Enumeration. Early greedy enumeration algorithms typically follow two strategies [38]. One is a top-down approach [21], such as the Drop algorithm [13], which iteratively removes candidate indexes with minimal impact on workload execution time until the budget constraint is met. The other is a bottom-up approach [14], exemplified by the Autoadmin algorithm proposed by Chaudhuri and Narasayya [12], which starts with an empty set and adds indexes until the maximum budget is reached. Recently, reinforcement learning-based methods, such as those by Lan et al. [11] and Sadri et al. [41], have shown greater adaptability and efficiency. SmartIX [42] introduces add/delete operations based on environmental state.
Workload Compression. To enhance the efficiency of index tuning, many approaches perform workload compression prior to tuning. Sadri et al. [41] identify templates based on the non-parameterized portions of query statements, retaining one query for each template. However, the query plans of a template can vary greatly depending on parameter differences, resulting in substantially different benefits calculated for the same index configuration and causing significant workload information loss. Furthermore, RIBE [39] clusters queries based on operators within the query plan, ensuring that only redundant information from the workload is eliminated. Besides these clustering-based methods, GSUM [2] and ISUM [43] define various query features and design greedy algorithms that flexibly determine the number of queries in the workload.
8 Conclusion
In this paper, we address a critical yet unexplored question: Can uncertainty enable better learning-based index tuning? To answer this question, we focus on the core component of index tuning, prioritizing improvements to the learning-based BE model. Our investigation is divided into two sub-questions: (1) Can UQ methods accurately measure the reliability of model results? (2) How can a BE model that provides uncertainty estimates be integrated into an index advisor, and what are the implications of this integration?
To address these two questions, we first analyze the origins and characteristics of model uncertainty in BE tasks. Unlike existing general UQ methods, we design a new hybrid quantification framework tailored to the task’s characteristics and propose an alternative version to enhance efficiency. Experimental validation shows that our method outperforms existing quantification methods in five out of six scenarios. The alternative version significantly improves UQ efficiency, albeit with a slight reduction in quantification accuracy. Next, we verified that the Beauty framework leverages the stability of what-if tools to supplement the learning-based BE model. In end-to-end experiments, this framework significantly improves the quality of index tuning results.
Our findings confirm that integrating uncertainty can enhance learning-based index tuning. Our framework focuses on the BE problem, marking an initial exploration in this area. Future research will investigate the role of uncertainty in other index advisor modules.
Acknowledgments
This work was partially supported by the National Natural Science Foundation of China (Grant No. 62072138).

### v2

@@
-Can Uncertainty Quantification Enable Better Learning-based Index Tuning?
+Can Uncertainty Quantification Improve Learned Index Benefit Estimation?
 Abstract
@@
 Index tuning, database, benefit estimation, uncertainty quantification.
-1 Introduction
+I Introduction
 Indexes are a crucial means for accelerating database queries.
@@
 (1) Poor stability: Learning-based models typically outperform what-if tools on data similar to their training sets, but they can generate highly erroneous results when query patterns shift or training data is insufficient.
-As shown in Figure 1, we compare a learning-based BE method with a what-if tool on varying proportions of out-of-domain (OOD) samples.
-When the proportion reaches around 25%, the mean errors of two methods are comparable, but the mean error of the learning-based method becomes larger than the what-if tool as the OOD sample proportion increases.
+As shown in Figure 1, we compare a learning-based BE method with what-if tools on varying proportions of out-of-domain (OOD) samples.
+When the proportion reaches around 25%, the mean errors of two methods are comparable, but the mean error of the learning-based method becomes larger than the what-if tools as the OOD sample proportion increases.
 Figure 1: Mean errors of a learning-based BE model and a what-if tool.
@@
 When the model’s estimate for a test sample shows high uncertainty, the result is more likely to significantly deviate from the true value.
-This technique, known as uncertainty quantification (UQ), has been widely studied in fields where model errors have serious consequences, such as medical image detection [8], autonomous driving [9], and drug analysis [10].
+This technique, known as uncertainty quantification (UQ), has been widely studied in fields where model errors have serious consequences, such as medical image detection, autonomous driving, and drug analysis [8, 9].
 By using this approach, we can return only the reliable benefit estimation results to the index advisor, while utilizing what-if tools to obtain the remaining results, thereby improving the reliability of the BE module.
@@
 (1) We conducted a systematic study of the sources of uncertainty in learning-based BE models within the context of index tuning.
-Based on the characteristics of these tasks, we proposed a novel hybrid UQ method that allows for a flexible trade-off between efficiency and accuracy, depending on the specific characteristics of the BE task (Section 4).
+Based on the characteristics of these tasks, we proposed a novel hybrid UQ method that allows for a flexible trade-off between efficiency and accuracy, depending on the specific characteristics of the BE task (Section IV).
 Our experiments demonstrate that the proposed UQ method outperforms general-purpose UQ methods in five out of six scenarios.
 (2) We introduced Beauty, the first uncertainty-aware BE framework, which combines the strengths of learning-based models with what-if tools, enabling uncertainty-driven model updates.
-Our experiments across six benchmarks with varying budgets show that the Beauty framework significantly improves index tuning performance (Section 6).
-(3) In our experiments, we applied existing UQ methods to various learning-based BE models, creating a total of 16 different models (Section 5).
+Our experiments across six benchmarks with varying budgets show that the Beauty framework significantly improves index tuning performance (Section VI).
+(3) In our experiments, we applied existing UQ methods to various learning-based BE models, creating a total of 16 different models (Section V).
 We thoroughly evaluated the impact of these methods on model accuracy and efficiency, confirming the superiority of our proposed UQ method.
-2 Background
-2.1 Index Advisors
+II Background
+II-A Index Advisors
 Index advisors are responsible for index tuning, that is, recommending indexes for databases [10].
@@
 is recommended for the database. Clearly, maximizing $B(W,\mathcal{I}_{0},\mathcal{I})$ is equivalent to minimizing the total cost $c(W,\mathcal{I})=\sum_{i=1}^{n}w_{i}\cdot c(q_{i},\mathcal{I})$.
-2.2 Learning-based Benefit Estimation (BE) Models
+II-B Learning-based Benefit Estimation (BE) Models
 The benefit estimator plays a crucial role in index tuning as its efficiency directly impacts the overall performance of index tuning, and its accuracy affects the behavior of the index configuration enumerator [14, 19, 20].
@@
 One approach such as DISTILL [1] involves training multiple lightweight BE models for each query template to collectively estimate benefits.
-The other approach is to train a single complex BE model using historical query execution data, either to completely replace the what-if tool or to adjust the outputs of the what-if tool.
+The other approach is to train a single complex BE model using historical query execution data, either to completely replace the what-if tools or to adjust the outputs of the what-if tool.
 The latter approach is the main focus of our study.
@@
 Predictor: The predictor, typically a MLP, takes the hidden vector $\mathbf{v}^{h}$ as input and generates the final output $\hat{y}=\text{MLP}(\mathbf{v}^{h})$ which represents an estimate of either $c(q,\mathcal{I})$ or $B(q,\mathcal{I}_{0},\mathcal{I})$.
-2.3 Uncertainty in Machine Learning (ML) Models
+II-C Uncertainty in Machine Learning (ML) Models
 Uncertainty is the quality or state of being uncertain, which arises in every stage of a ML task, including data acquisition, feature encoding, model design, model training, and model inference [8].
-Naturally, uncertainty is inherent in ML-based index tuning, which will be explained in subsection 4.2.
+Naturally, uncertainty is inherent in ML-based index tuning, which will be explained in subsection IV-B.
 As illustrated in Figure 4, there are two different types of uncertainty, often referred to as aleatoric uncertainty and epistemic uncertainty [9].
@@
 Uncertainty Quantification (UQ): A number of model-independent and model-dependent methods have been proposed to quantify the two types of uncertainty.
-The first category includes Bayesian neural networks (BNNs) [25, 26], Monte Carlo dropout (MCD) methods [27], and ensemble methods [28, 29], which do not rely on the specific models in which the uncertainty is quantified.
+The first category includes Bayesian neural networks (BNNs) [22, 23], Monte Carlo dropout (MCD) methods [24], and ensemble methods [25], which do not rely on the specific models in which the uncertainty is quantified.
 The second category incorporates UQ into model-specific structures.
@@
 In this paper, we will reproduce these methods in the experimental evaluation and compare them with our proposed hybrid UQ approach.
-3 Problem
+III Problem
 Since collecting training data is costly, ML models inevitably encounter out-of-distribution (OOD) data, leading to unreliable outputs.
 The unreliable results produced by learning-based benefit estimators can lead index advisors to select inappropriate indexes, significantly degrading database performance.
-Therefore, as discussed in section 1, we will explore how UQ techniques can be leveraged to determine whether the results produced by learning-based BE models are reliable and how to integrate uncertainty-aware benefit estimators into index advisors.
+Therefore, as discussed in section I, we will explore how UQ techniques can be leveraged to determine whether the results produced by learning-based BE models are reliable and how to integrate uncertainty-aware benefit estimators into index advisors.
 Specifically, this paper addresses the following three key research questions:
@@
 In this paper, the first problem is how to precisely quantify the uncertainty in a learning-based BE model.
-Rather than applying the existing general-purpose UQ methods directly, we propose a hybrid quantification method in section 4 tailored to the unique characteristics of the BE problem, providing more precise quantification of uncertainty than the general-purpose methods.
+Rather than applying the existing general-purpose UQ methods directly, we propose a hybrid quantification method in section IV tailored to the unique characteristics of the BE problem, providing more precise quantification of uncertainty than the general-purpose methods.
 Figure 5: Joint distribution of model error and model uncertainty evaluated using UQ methods BNN [27], MCD [24], HSTO [26], and Ensemble [25]. $r$: Pearson’s correlation coefficient.
 Q2: What impacts does UQ have on a BE model?
-In section 5, we will conduct an empirical study to understand the impacts of UQ on BE models by evaluating the accuracy, training cost, and inference cost of the BE models enhanced with UQ.
-In subsection 4.4, we will also explore a new method for determining when a BE model needs to be updated based on quantified uncertainty.
+In section V, we will conduct an empirical study to understand the impacts of UQ on BE models by evaluating the accuracy, training cost, and inference cost of the BE models enhanced with UQ.
+In subsection IV-D, we will also explore a new method for determining when a BE model needs to be updated based on quantified uncertainty.
 Q3: What impacts does a UQ-enhanced BE model have on the index advisor?
-In section 6, we will evaluate 8 UQ-enhanced BE models with 6 workloads across 36 scenarios to investigate the effects of the accuracy and efficiency of UQ on the index advisor.
+In section VI, we will evaluate 8 UQ-enhanced BE models with 6 workloads across 36 scenarios to investigate the effects of the accuracy and efficiency of UQ on the index advisor.
 Due to space constraints, we focus exclusively on BE models based on neural networks (NNs).
 NN-based models have been receiving increasing attention and generally achieve higher accuracy.
-The approach proposed in this paper can be easily extended to non-NN models by replacing the UQ methods for non-NN models such as [33, 34].
-4 The Beauty Framework
-To answer the questions in section 3, we propose our new framework for benefit estimation called Beauty (Benefit estimation with uncertainty).
-4.1 Overview
-The process of index tuning with Beauty is illustrated in Figure 6.
-After generating an index configuration by the enumerator, the advisor uses the uncertainty-aware learning-based BE model to estimate both the benefits that the configuration brings to the queries in the workload and the confidence levels of the estimated benefits.
-Then, the result filter examines the estimated benefits and their uncertainty with respect to a specified uncertainty threshold.
-The estimated benefits with uncertainty below the threshold are deemed reliable and are returned to the enumerator.
-For the remaining estimated benefits on which the BE model has low confidence due to insufficient training data or feature extraction deficiencies, the what-if tool that is usually more robust than learning-based BE models are invoked instead for benefit estimation.
-Figure 6: The index tuning process with Beauty.
-4.2 Uncertainty in Learning-based BE Models
-The existing UQ methods are designed for general ML models and scenarios, lacking specific knowledge necessary for benefit estimation.
+The approach proposed in this paper can be easily extended to non-NN models by replacing the UQ methods for non-NN models such as [28].
+IV The Beauty Framework
+To answer the questions in section III, we propose our new framework for benefit estimation called Beauty (Benefit Estimation with UncertainTY).
+IV-A Overview
+The index tuning process with Beauty is illustrated in Figure 6.
+After the enumerator generates an index configuration, the advisor uses the uncertainty-aware learning-based BE model to estimate both the benefits that the index configuration brings to the queries in the workload and the uncertainties of the estimated benefits.
+Then, the result filter examines the estimated benefits and their uncertainties with respect to a specified uncertainty threshold.
+The estimated benefits with uncertainties below the threshold are deemed reliable and returned to the enumerator.
+For the remaining estimated benefits with uncertainties above the threshold due to insufficient training data or deficiencies in feature extraction, the what-if tool is invoked for benefit estimation because it is usually more robust than learning-based BE models.
+Figure 6: Index tuning process with Beauty.
+IV-B Uncertainties in Learning-based BE Models
+The existing UQ methods are designed for general ML models and scenarios, lacking specific knowledge necessary for BE.
 To overcome this limitation, we propose a hybrid UQ method tailored to the unique characteristics of the BE problem, providing more precise quantification of uncertainty than the general-purpose methods.
-To this end, we conduct the first analysis of the sources of uncertainty in the BE problem and their relationships to the structures of learning-based BE models, which lays foundations for designing the uncertainty-aware BE model in Beauty.
-Uncertainty emerges throughout the whole process of learning-based BE models from design to training and to inference. We divide the uncertainty into five categories which are summarized in Table I.
+To this end, we conducted the first analysis of the sources of uncertainties in the BE problem and their relationships with the structures of learning-based BE models, which lays the foundations for designing the uncertainty-aware BE model in Beauty.
+Uncertainties emerge throughout the entire process of learning-based BE models, from model design to model training, and to model inference.
+Uncertainties can be divided into five categories as summarized in Table I.
 TABLE I: Five categories of uncertainty in learning-based BE models.
@@
 Estimating the benefit of an index configuration for a query is a complex task, as it requires invoking the query optimizer to retrieve the plans of the query before and after creating the indexes in the configuration.
-Existing learning-based estimators only model the information most relevant to this query such as the index configuration, the current query plan, and the plan chosen by the optimizer after virtually creating the indexes in the configuration.
+Existing learning-based estimators only model the information most relevant to this query, such as the index configuration, the current query plan, and the plan chosen by the optimizer after virtually creating the indexes in the configuration.
 However, some overlooked factors can also significantly affect query execution costs.
-One such factor is data distributions which determine the cardinalities of intermediate results produced during the execution of a query.
+One such factor is data distributions that determine the cardinalities of intermediate results produced during the execution of a query.
 Existing learning-based estimators either do not model data distributions or model them in an over-simplified manner.
 Cardinality estimation itself is challenging and is a key research focus in the AI4DB field.
-State-of-the-art data-driven cardinality estimation methods are often built on deep-learning models which are more complex than those used in the existing BE models.
+State-of-the-art data-driven cardinality estimation methods are often built on deep learning models that are more complex than those used in existing BE models.
 Another factor that significantly affects query execution costs is the configuration of the database system such as the knobs work_mem and shared_buffers.
-Finally, the implementation of the query optimizer itself can alter the selected plan for a query and varies its execution cost.
-Currently, none of the learning-based estimators has taken all these relevant information into account.
-The ignorance of these information leads to epistemic uncertainty.
+Finally, the implementation of the query optimizer itself can alter the selected plan for a query and vary its execution cost.
+Currently, none of the learning-based estimators has taken all this relevant information into account.
+The ignorance of related information leads to epistemic uncertainty.
 U2: Epistemic uncertainty caused by flawed feature extraction.
@@
 Therefore, if a BE model estimates benefits based on observed query execution time, the model’s outputs must exhibit randomness, and the randomness is related to the characteristics of queries, resulting in heteroscedastic aleatoric uncertainty.
-4.3 Hybrid Quantification Method
+IV-C Hybrid Quantification Method
 In this section, we propose a hybrid method for quantifying uncertainty in a learning-based BE model.
 We first describe which modules in the BE model is each category of uncertainty mainly quantified from, followed by the design of specific quantification methods.
-4.3.1 Modules for Quantifying Uncertainty
-As described in subsection 2.2, there are three main modules in a learning-based BE model: the feature extractor, the hidden vector encoder, and the predictor.
-Here, we analyze the relationship between the five categories of uncertainty listed in subsection 4.2 and the modules of the BE model, which lays the foundation for our hybrid UQ method.
+IV-C1 Modules for Quantifying Uncertainty
+As described in subsection II-B, there are three main modules in a learning-based BE model: the feature extractor, the hidden vector encoder, and the predictor.
+Here, we analyze the relationship between the five categories of uncertainty listed in subsection IV-B and the modules of the BE model, which lays the foundation for our hybrid UQ method.
 Quantifying Uncertainty from Hidden Vector Encoder Module.
@@
 Thus, uncertainty U1 and U2 can also be quantified within the predictor module.
-4.3.2 Quantifying Uncertainty from Hidden Vector Encoder
+IV-C2 Quantifying Uncertainty from Hidden Vector Encoder
 Given a learning-based BE model, uncertainties U3 and U4 represent the model’s fitting error with respect to the benefit function $f$, which is reflected in the quality of the hidden feature vector $\mathbf{v}^{h}$ extracted by the hidden vector encoder module.
@@
 The rationale behind this approach lies in the fact that for the decoder to accurately reconstruct $V^{1}$ from $V^{h}$, two essential conditions must be satisfied.
-First, the structure of the hidden vector encoder must be appropriate.
-The network structure of the decoder should be symmetric to that of the encoder.
+First, the mechanism and architecture implemented by the hidden vector encoder must be well-designed.
+The decoder typically employs identical mechanisms with a mirrored network structure.
 Only if the encoder’s structure can gradually eliminate unimportant features and reduce dimensionality, the decoder can reverse this process and reconstruct $V^{1}$ based on the essential features.
@@
 In engineering practice, MSE can be replaced by weighted MSE, and the practitioners can properly set weights to the dimensions of $\mathbf{v}^{1}_{i}$ and $\hat{\mathbf{v}}^{1}_{i}$.
-4.3.3 Quantifying Uncertainty from Predictor
+IV-C3 Quantifying Uncertainty from Predictor
 After using the decoder to quantify uncertainty U3 and U4 in the hidden vector encoder, we proceed to quantify uncertainties U1, U2, and U5 in the predictor module.
@@
 $p(\mathbf{X}_{W}|D)$ $=\frac{p(D|\mathbf{X}_{W})p(\mathbf{X}_{W})}{p(D)}$
-[…]
+$=\frac{p(D|\mathbf{X}_{W})p(\mathbf{X}_{W})}{\idotsint p(D|\mathbf{X}_{W})p(\mathbf{X}_{W})d\mathbf{X}_{W}}.$
 The prior distribution $p(\mathbf{X}_{W})$ should reflect prior beliefs about the distribution of the weights $W$ before observing any data.
-Solving this equation directly can be challenging, so variational inference (VI) is often employed [36, 26].
+Solving this equation directly can be challenging, so variational inference (VI) is often employed [23].
 Blundell et al.
@@
 Dropout layers are commonly used in deep NNs to prevent overfitting.
-During training, a dropout layer randomly sets elements of the input tensor to 0 with probability $p\in[0,1]$ and scales up the remaining elements by $1/(1-p)$ to maintain the sum over all inputs.
+During training, a dropout layer randomly sets elements of the input tensor to 0 with probability $p\in[0,1]$ and scales up the remaining elements by 1/(1p)1/(1-p) to maintain the sum over all inputs.
 The zeroed elements are chosen independently for each forward pass.
@@
 $\mathcal{U}_{2}=\textrm{Var}(\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{m}).$ (3)
-4.3.4 Result Filter
+IV-C4 Result Filter
 Once the BE model with UQ capability (referred to as the BE-UQ model) provides the estimated benefit $\hat{y}$ and the quantified uncertainties $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, the Result Filter module (as shown in Figure 6) determines whether $\hat{y}$ should be returned to the Estimator.
@@
 Specifically, the returned result $\hat{B}(q,\mathcal{I}_{0},\mathcal{I})$ should follow:
-$\hat{B}(q,\mathcal{I}_{0},\mathcal{I})=\begin{cases}\hat{y}&\text{if }\mathcal{U}_{1}\leq\theta_{1}\ \text{and}\ \mathcal{U}_{2}\leq\theta_{2},\\ \text{What-if tools}&\text{otherwise}.\end{cases}$
-The domain of $\theta_{1}$ and $\theta_{2}$ is $[0,+\infty)$, and their values vary based on factors such as model structure, hyperparameters, query workload, and the database instance.
-Setting fixed, universal thresholds for $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ is impractical.
-Therefore, we adopt a context-dependent approach to define relative thresholds $\theta_{1}$ and $\theta_{2}$: for a new sample, if the model’s uncertainty regarding it is comparable to the uncertainty observed during training when the model performed optimally with minimal uncertainty, the result is deemed reliable.
-Specifically, after training the BE-UQ model, we perform inference for each training sample $(q,\mathcal{I}_{0},\mathcal{I})$ and gather the quantified uncertainties $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$.
-The 90th percentile of $\mathcal{U}_{1}$ (or $\mathcal{U}_{2}$) across all training samples is designated as $\theta_{1}$ (or $\theta_{2}$).
-4.4 Model Deployment
-4.4.1 Model Training
-Compared to the original BE model, our BE-UQ model incorporates two additional modules: the decoder and MCD.
-Since the decoder cannot be trained concurrently with the original BE model, a redesigned training process is required.
-In our index advisor, the BE-UQ model is trained in two distinct phases.
-First, the BE model—which includes the feature extractor, hidden vector encoder, and Predictor—is trained following the original BE model’s method.
-In the second phase, the parameters of the feature extractor and hidden vector encoder are frozen, and the decoder is trained by minimizing the MSE loss.
-The rationale behind this training method is that BE remains the primary objective of the BE-UQ model, while UQ is secondary.
-Our approach ensures that the accuracy of the BE model is maintained, regardless of whether UQ is utilized.
-By preserving the accuracy of the hidden vector encoder, we ensure that the input information is effectively encoded in the intermediate representation $V^{h}$.
-This, in turn, enables the decoder to accurately reconstruct the input vector set $\hat{V}^{1}$ and correctly quantify uncertainty.
-Although there are alternative ways to train the BE-UQ model, such as jointly training the hidden vector encoder and decoder followed by training the Predictor, this method shifts the encoder’s focus toward UQ rather than BE, which conflicts with our objectives.
-Therefore, our proposed training method is better suited for the index tuning task.
-4.4.2 Uncertainty-Driven Model Updating
-The performance of the BE-UQ model may decline when the database or query workload changes.
-To maintain the accuracy and stability of the model, it must be updated promptly when performance starts to decrease.
-There are two common ways to determine whether a machine learning model requires updating.
-The first category samples predictions during model inference and compares them with actual values.
-A model update is needed if the error is significantly large.
-However, this approach has two main drawbacks.
-First, the actual benefits $B(q,\mathcal{I}_{0},\mathcal{I})$ for the sampled instances $(q,\mathcal{I}_{0},\mathcal{I})$ must be known.
-To determine $B(q,\mathcal{I}_{0},\mathcal{I})$, the indexes in $\mathcal{I}$ must be created, and the query $q$ must be executed based on those indexes, which incurs substantial time and space costs.
-Second, this method may cause delayed updates due to sampling errors.
-The second category triggers model updates when a decrease in system efficiency is detected.
-However, this approach causes updates to lag behind the onset of performance regression.
-More importantly, in the context of index tuning, the optimal index is unknown.
-Even if an erroneous BE result leads to the creation of a suboptimal index, system performance may still improve slightly.
-Thus, it is difficult to determine whether the model requires updating based on this result.
-In the Beauty framework, we have a more accurate and timely metric to determine whether an update is needed: if the what-if tool estimates benefits for a significant portion of test samples $(q,\mathcal{I}_{0},\mathcal{I})$, this likely indicates that the BE-UQ model no longer fits the current dataset, workload, or system conditions and requires updating.
+$\hat{B}(q,\mathcal{I}_{0},\mathcal{I})=\begin{cases}\hat{y_{1}}&\text{if }\mathcal{U}_{1}\leq\theta_{1}\ \text{and}\ \mathcal{U}_{2}\leq\theta_{2},\\ \hat{y_{2}}&\text{otherwise},\end{cases}$
+where $\hat{y_{1}}$ and $\hat{y_{2}}$ represent the results obtained from the learning-based model and the what-if tool, respectively.
+The domain of $\theta_{1}$ and $\theta_{2}$ is $0,+\infty)$, and their values are influenced by factors such as model architecture, hyperparameters, query workload, and the database instance.
+Consequently, it is impractical to set fixed, universal thresholds for $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$.
+However, we observe two key facts: (1) A well-trained learning-based model yields accurate and reliable results on the training dataset, where uncertainty is minimized.
+(2) The uncertainties quantified for the training instances typically follow a long-tailed distribution, with over 94% of cases in the 54 experimental results presented in subsection V-C conforming to this trend.
+Based on these observations, we propose a context-aware method to determine $\theta_{1}$ and $\theta_{2}$.
+After training the BE-UQ model, we infer the BE-UQ model for each training sample $(q,\mathcal{I}_{0},\mathcal{I})$ and collect the uncertainties $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ quantified for $(q,\mathcal{I}_{0},\mathcal{I})$.
+A naïve method is to set $\theta_{i}=\max(\mathcal{U}_{i})$.
+However, due to the long-tailed distribution of $\mathcal{U}_{i}$, the presence of a few large uncertainties (outliers) can make $\theta_{i}$ unnecessarily large, causing many unreliable samples not identified.
+Therefore, we refer to the Interquartile Range (IQR) method [30] to set
+$\theta_{i}=\min(P_{75}(\mathcal{U}_{i})+\alpha\times(P_{75}(\mathcal{U}_{i})-P_{25}(\mathcal{U}_{i})),\max(\mathcal{U}_{i})),$
+where $P_{25}(\mathcal{U}_{i})$ and $P_{75}(\mathcal{U}_{i})$ are the 25th and 75th percentiles of the collected values of $\mathcal{U}_{i}$, respectively.
+The parameter $\alpha$ is typically set within the range of $1$ to $1.5$.
+Due to the long-tail distribution of $\mathcal{U}_{i}$, we usually have $\theta_{i}\ll\max(\mathcal{U}_{i})$, which can significantly increase the recall of unreliable samples.
+IV-D Model Deployment
+IV-D1 Model Training
+Compared to the original BE model, our BE-UQ model incorporates two additional modules: the decoder and the MCD.
+Since the decoder cannot be trained concurrently with the original BE model, the training process is redesigned.
+In particular, the BE-UQ model is trained in two distinct phases.
+In phase 1, the BE model—which includes the feature extractor, hidden vector encoder, and Predictor—is trained using the original method proposed for the BE model.
+In phase 2, the parameters of the feature extractor and the hidden vector encoder are frozen, and the decoder is trained to minimize the MSE loss.
+The framework and hyperparameters for the decoder can be selected independently of those for the encoder.
+The rationale behind this training method is that BE is the primary objective of the BE-UQ model, while UQ is secondary.
+Our training method ensures the accuracy of BE, regardless of whether UQ is enabled.
+By preserving the accuracy of the hidden vector encoder, we guarantee that the input information is effectively encoded in the intermediate representation $V^{h}$.
+This, in turn, enables the decoder to accurately reconstruct the input vector set $\hat{V}^{1}$ and correctly quantify uncertainties.
+IV-D2 Uncertainty-Driven Model Updating
+The performance of the BE model may degrade when the underlying database or query workload undergoes changes.
+To maintain the model’s accuracy and stability, it is imperative to promptly initiate updates when a decline in performance is detected.
+There are two primary approaches to determining when a model requires an update.
+The first approach involves sampling predictions during model inference and comparing them to actual values.
+An update is triggered if the error exceeds a predefined threshold.
+However, this method necessitates the creation of indexes in $\mathcal{I}$ to obtain the actual benefits $B(q,\mathcal{I}_{0},\mathcal{I})$, thereby incurring substantial time and space costs.
+Additionally, this approach may result in delayed updates due to sampling errors.
+The second approach triggers model updates upon detection of a decrease in system efficiency.
+However, this method suffers from a significant drawback: it causes updates to lag behind the onset of performance regression.
+In the Beauty framework, we have a more accurate and timely metric to determine whether an update is needed: if the what-if tools estimate benefits for a significant portion of test samples $(q,\mathcal{I}_{0},\mathcal{I})$, this likely indicates that the BE-UQ model no longer fits the current dataset, workload, or system conditions and requires updating.
 Additionally, the uncertainties $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ returned by the model can help guide the update process.
-(1) If the what-if tool estimates the benefits for a significant portion of test samples $(q,\mathcal{I}_{0},\mathcal{I})$ due to $\mathcal{U}_{1}>\theta_{1}$, this suggests that the BE-UQ model cannot provide effective latent vectors under the current workload.
+(1) If the what-if tools estimate the benefits for a significant portion of test samples $(q,\mathcal{I}_{0},\mathcal{I})$ due to $\mathcal{U}_{1}>\theta_{1}$, this suggests that the BE-UQ model cannot provide effective latent vectors under the current workload.
 For example, if there have been substantial changes in query patterns or index configurations, the BE-UQ model may fail to produce accurate results.
 In this case, more data must be collected from the current workload to retrain or fine-tune the model.
-(2) If the what-if tool estimates the benefits for a significant portion of test instances $(q,\mathcal{I}_{0},\mathcal{I})$ due to $\mathcal{U}_{2}>\theta_{2}$, this indicates that the BE-UQ model is unable to distinguish between too many queries and index configurations in the current workload.
+(2) If the what-if tools estimate the benefits for a significant portion of test instances $(q,\mathcal{I}_{0},\mathcal{I})$ due to $\mathcal{U}_{2}>\theta_{2}$, this indicates that the BE-UQ model is unable to distinguish between too many queries and index configurations in the current workload.
 Due to insufficient modeling information or flaws in feature extraction, the model estimates a wide range of benefits for the same test instance, leading to unreliable predictions.
 In such cases, the designer must identify the missing information or flaws in feature extraction and redesign the model based on the current workload.
-4.4.3 Improving Efficiency.
+IV-D3 Efficient Version.
 During index tuning, the efficiency of benefit estimation significantly affects the tuning time and, in some cases, can influence the final tuning quality, particularly with algorithms like DTA that can be interrupted at any point.
@@
 Consequently, the trained model will inevitably encounter workloads that deviate from those seen during training.
-Therefore, when the feature extraction module is effective and the encoded information is sufficient, $\mathcal{U}_{2}$ quantification can be disabled.
-While this may slightly reduce the accuracy of the UQ, it significantly improves the inference efficiency of the BE-UQ model.
-5 Empirical Study on Uncertainty Quantification
-In this section, we compare different UQ methods from three aspects, namely the UQ capability, the efficiency and the impact on the accuracy of BE.
-5.1 Datasets Construction
-Three famous database benchmarks are adopted in the empirical study. Table II summarizes the databases and the workloads.
-TPC-H: Following Leis et al.
-[38, 39], we randomly generated 30 queries from each template in TPC-H, excluding templates 2, 17, and 20.
-These were excluded because their queries have significantly longer execution times, which would disproportionately affect the overall workload cost.
-Indexes accelerating these queries are typically preferred over those for other templates.
-TPC-DS: We also randomly generated 30 queries from each template in TPC-DS. Templates 4, 6, 9, 10, 11, 32, 35, 41, and 95 were excluded for the same reason given above.
-JOB: This workload consists of 30 queries randomly generated from each template in JOB.
-To increase the complexity of index tuning, we ingested synthetic queries into these workloads using the method devised by Yu et al.
-[39].
-The enhanced workloads are referred to as TPC-H+, TPC-DS+ and JOB+, respectively.
+Therefore, in scenarios where high efficiency is required, the quantification of $\mathcal{U}_{2}$ can be disabled.
+As will be shown in the evaluation in section V, eliminating the quantification of $\mathcal{U}_{2}$ can significantly increase the efficiency of model inference, while slightly decreasing the accuracy of UQ.
+V Empirical Study on Uncertainty Quantification
+In this section, we integrate UQ methods into the BE model, thereby establishing BE-UQ models.
+We will systematically evaluate each BE-UQ model across three dimensions: uncertainty quantification capability, benefit estimation accuracy, and inference efficiency.
 TABLE II: Summary of databases and workloads and the maximum runtime for index tuning algorithm DTA.
-Benchmark DB SF/Size # Queries # Templates # Synthetic Queries Avg. # Plan Nodes Maximum Runtime (mins)
-TPC-H SF = 10 570 19 0 14 20
-TPC-DS SF = 1 2700 90 0 29 45
-JOB 13GB 3390 113 0 24 60
-TPC-H+ SF = 10 950 19 380 10 30
-TPC-DS+ SF = 1 4500 90 1800 19 60
-JOB+ 13GB 5650 113 2260 16 75
-To the best of our knowledge, no systematic experiments have previously explored uncertainty involved in index tuning tasks.
-To investigate the impacts of UQ on the accuracy and the efficiency of learning-based benefit estimators, as well as their ability to quantify uncertainty, we designed the following experimental methodology:
-We ran the Autoadmin and Extend algorithms on three distinct workloads: TPC-H+, TPC-DS+, and JOB+, under various budget constraints to collect index configurations for tuning.
-After materializing these configurations, we gathered query execution data, represented as [(query, index config) $\rightarrow$ benefit], and formed a comprehensive dataset.
-This dataset was then divided into three segments based on query templates and indexes:
-Training Dataset: Used for model training, containing approximately 50% of the total data. This subset only includes some of the query templates and indexes.
-Test Dataset:Comprises about 15% of the total data, where each query or index configuration appeared in the training set but was not simultaneously seen by the model.
-This segment tests the model’s generalization capability within in-domain data.
-Uncertainty Dataset:This dataset accounts for about 35% of the total data.
-It includes queries and indexes that are not present in the first two datasets, and no similar queries (from the same template) appear.
-For each model mentioned in Subsubsection 5.2.2, we will test using these samples, excluding those with errors below the 90th percentile of the Training Dataset errors.
-This way, the remaining samples are those the model fails to predict correctly due to U1-U4, and the model should assign a high uncertainty value to them.These samples are designated as Uncertainty Samples.
-There are two points to note.
-First, the samples that different models fail to predict accurately vary, resulting in an inconsistent number of Uncertainty Samples, but the difference does not exceed 10%.
-Second, we use query costs to calculate the benefit rather than the execution time of the query, so we do not include U5-type uncertainty samples here.
-5.2 Experimental Setup
-5.2.1 BE Models
-In this experiment, we adopt two well-known learning-based BE models:
+Benchmark DB SF/Size # Queries # Templates
+# Synthetic
+Queries
+Avg. #
+Plan Nodes
+Maximum
+Runtime
+TPC-H+ SF = 10 950 19 380 10 90
+TPC-DS+ SF = 10 4500 90 1800 19 120
+JOB+ 13GB 5650 113 2260 16 180
+V-A Construction of Datasets
+We prepared databases and query workloads based on three famous benchmarks, TPC-H, TPC-DS, and JOB.
+For each benchmark, we create a workload that consists of two groups of queries.
+(1) For each query template provided in the benchmark, we randomly generated $30$ queries.
+Following the procedure commonly conducted in previous studies [32, 17, 33], we exclude templates 2, 17, and 20 from TPC-H and exclude templates 4, 6, 9, 10, 11, 32, 35, 41, and 95 from TPC-DS.
+These templates were excluded because queries generated using these templates have much longer execution times than queries generated using other templates.
+If these templates were used, index tuning tends to recommend indexes that only accelerate queries generated using these templates.
+(2) To increase the difficulty of benefit estimation and index tuning, complex synthetic queries that include joins or nested queries were randomly generated.
+To ensure each synthetic query to yield at least one result, we adopted the method [33] in which the results are generated first, followed by the parameters in the query.
+We call the extended benchmarks with synthetic queries TPC-H+, TPC-DS+, and JOB+, respectively.
+Their statistics are given in Table II.
+To the best of our knowledge, there have been no comprehensive experimental studies on the performance of UQ methods integrated with learning-based BE models in the context of index tuning.
+Therefore, we design the following experiments.
+First, for each benchmark, we construct a comprehensive dataset $D$ that consists of records of the form (query, indexes, benefit).
+In particular, we use the algorithms AutoAdmin [11] and Extend [19] to generate candidate index configurations for the query workload.
+For each generated index configuration $\mathcal{I}$, we run each query $q$ in the workload before and after creating the indexes in $\mathcal{I}$ and record the actual benefit $b$ that these indexes bring to $q$, thus obtaining a record $(q,\mathcal{I},b)$ in $D$.
+To evaluate the performance of UQ methods integrated with learning-based BE models, we divide $D$ into $3$ disjoint partitions $D_{train}$, $D_{test}$ and $D_{eval}$, where $D_{train}$ is used to train the BE-UQ model, $D_{test}$ is used to test the BE accuracy of the BE-UQ model, and $D_{eval}$ is used to evaluate the UQ capability of the BE-UQ model.
+Notably, we deliberately select $D_{eval}$ to exhibit substantially distinct distributional characteristics compared to $D_{train}$ and $D_{test}$.
+Due to the distinction in the distributions of $D_{train}$ and $D_{eval}$, the BE-UQ model trained on $D_{train}$ is supposed to give high uncertainties for the samples in $D_{eval}$.
+Therefore, we can use the fraction of samples in $D_{eval}$ that are identified as unreliable by the BE-UQ model using the method proposed in subsubsection IV-C4 as a metric of UQ performance.
+In particular, $D_{eval}$ is constructed as follows.
+First, randomly select some query templates and columns from the benchmark.
+Then, pick the records $(q,\mathcal{I},b)$ in $D$ into $D_{eval}$ if $q$ is generated from a selected template, or $\mathcal{I}$ contains the selected columns.
+We carefully select templates and columns so that $D_{train}$, $D_{test}$ and $D_{eval}$ approximately contain $50\%$, $15\%$, and $35\%$ of records in $D$, respectively.
+V-B Experimental Setup
+V-B1 BE Models
+In this experiment, we adopt two well-known learning-based BE models.
 (1) LIB.
-The BE model proposed in [7] encodes the operators and information in a query plan that can be influenced by indexes, forming a variable-length Index Optimizable Operations Set.
-After encoding through Multi-head Attention, the pooling operation retrieves $v^{h}$, culminating in the output of the BE via a MLP.
+The LIB model [7] encodes the operators and information in a query plan that can be influenced by indexes, forming a variable-length Index Optimizable Operations Set.
+After encoding through the Multi-head Attention mechanism, the pooling operation retrieves $v^{h}$, culminating in the output of BE via an MLP.
 (2) AMA.
-The model in [6] utilizes multiple channels to extract information and employs a weighted sum to compare query plans, selecting the faster option.
-Following [7], we replaced AMA’s classification layer with an MLP for benefit estimation.
-It is important to note that while we aim to maintain consistency with the original paper regarding the two BE models, complete consistency may not be achievable.
-Our primary focus is on studying the impact of the UQ method rather than directly comparing the two BE models.
-5.2.2 UQ Methods
+The AMA model [6] utilizes multiple channels to extract information from the plans provided by what-if tools and employs a weighted sum to compare these plans, ultimately selecting the faster option.
+Following [7], we replaced AMA’s classification layer with an MLP for BE.
+V-B2 UQ Methods
 We integrate five types of UQ methods into the aforementioned BE models, forming sixteen distinct BE-UQ models for our experiment:
@@
 This model integrates the hybrid quantification method proposed in this paper, resulting in LIB-Hybrid and AMA-Hybrid.
-We also compare an alternative version that utilizes efficiency enhancements, represented as LIB-AE and AMA-AE.
+Furthermore, we compare the efficiency-enhanced variants outlined in  subsubsection IV-D3, denoted as LIB-AE and AMA-AE.
 (2) BNN. We replace the MLP in the Prediction Module of both models with BNN, forming LIB-BNN and AMA-BNN.
@@
 The second variant applies MCD only to the last layer of the model’s Prediction Module, forming LIB-MCD-last and AMA-MCD-last.
-The third variant adopts the approach proposed in [37], which utilizes determinantal point processes for sampling to enhance diversity with minimal inference time increase, forming LIB-MCD-dpp and AMA-MCD-dpp.
+The third variant adopts the approach proposed in [31], which utilizes determinantal point processes for sampling to enhance diversity, forming LIB-MCD-dpp and AMA-MCD-dpp.
 In addition to the above methods, we also compare two other types of UQ methods.
@@
 Since this method relies on the Self-Attention module within the model, it is applicable only to LIB. We adopt the two versions mentioned in the original paper, resulting in LIB-STO and LIB-HSTO.
-5.2.3 Implementation
-We set the relative benefit $rb=\frac{B(q,\mathcal{I}_{0},\mathcal{I})}{c(q,\mathcal{I}_{0})}$ as the prediction target.
-To maintain compatibility with the LIB framework, we exclude data entries that exhibit negative benefits from our dataset.
+V-B3 Implementation
+We define the relative benefit as the prediction target:
+$\frac{B(q,\mathcal{I}_{0},\mathcal{I})}{c(q,\mathcal{I}_{0})}.$ (4)
+To maintain compatibility with the LIB framework, we exclude data entries that exhibit negative benefits from the training dataset.
 For each model discussed, we generate thirty sets of hyperparameters and train on the training dataset.
 Subsequently, we retain the model that demonstrates the best performance based on evaluations using the test dataset.
+Initially, the BE-UQ models constructed using our method implemented a mirror-symmetric decoder architecture with respect to their corresponding encoders, adhering to established conventions in this domain.
+However, when this symmetric configuration failed to achieve satisfactory convergence in the decoder component, we preserved the fundamental mechanisms (e.g., attention) while restructuring the neural architecture.
+Specifically, we systematically varied both the neuron density per layer and the network depth, thereby generating five distinct decoder variants.
+Following comprehensive retraining of each decoder configuration, we retained exclusively the variant that exhibited superior performance metrics.
 When quantifying uncertainty, we use the variance of results from five models as the quantification of uncertainty for the ensemble-based model.
 For other models, we perform inference $20$ times [31] for each batch sample, using variance as the quantification of uncertainty.
-(a) TPC-H+
-(b) TPC-DS+
-(c) JOB+
-Figure 8: Uncertainty distribution of models using BNN, MCD, Ensemble, and HSTO.
-Figure 9: Uncertainty distribution of models using Hybrid.
-Figure 10: Uncertainty distribution of models using AE.
-5.2.4 Hardware & Software
-The experiments were carried out on a Ubuntu server with two Intel Xeon 4210R CPUs (10 cores, 2.40GHz), 256GB of main memory, and one NVIDIA GeForce RTX 3060 GPU. The DBMS is PostgreSQL 12.13.
-5.3 Results on Uncertainty Quantification
-We present Raincloud plots illustrating the uncertainty distributions for sixteen models across various benchmarks.
-Figure 8 displays the plots for models based on BNN, MCD, Ensemble, and HSTO.
-Figures 9 and 10 show the plots for models based on Hybrid and AE, respectively.
-Each subplot depicts the uncertainty distribution of samples as calculated by the model for the training dataset, test dataset, and Uncertainty Samples.
-Consequently, the greater the uncertainty of the orange samples representing Uncertainty Samples in the graph and the higher their differentiation from other samples, the more precisely the model’s uncertainty modeling is demonstrated, enabling it to effectively distinguish uncertainty data.
-The red lines in the graphs represent the average uncertainties across the three datasets.
-TABLE III: Percentage of Uncertainty Samples exceeding the uncertainty threshold set by the Training Dataset.
-LIB-based Model TPC-H+ TPC-DS+ JOB+ AMA-based Model TPC-H+ TPC-DS+ JOB+
-BNN 13.03 40.20 16.04 BNN 45.26 65.32 53.67
-MCD-all 68.08 40.98 26.22 MCD-all 51.73 54.17 45.61
-MCD-last 61.79 29.32 5.92 MCD-last 21.76 21.27 22.11
-MCD-dpp 16.02 36.91 8.35 MCD-dpp 78.28 34.83 30.76
-Ensemble 79.01 53.44 29.32 Ensemble 38.73 77.62 60.22
-AE 71.85 31.13 30.93 AE 92.68 84.19 61.00
-Hybrid 81.80 57.65 32.88 Hybrid 91.74 78.13 47.82
-$\mathcal{U}_{1}$ 51.88 27.21 20.68 $\mathcal{U}_{1}$ 88.87 54.21 20.00
-$\mathcal{U}_{2}$ 59.70 43.46 20.76 $\mathcal{U}_{2}$ 54.93 47.50 34.87
-STO 68.41 34.86 23.51
-HSTO 63.64 35.13 24.58
-Table III further illustrates the filtering capabilities of various methods for Uncertainty Samples.
-Since Hybrid includes two UQ modules, we adopt the following approach to ensure comparability with other models: for $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, we select the 90th percentile of the uncertainty distribution from the Training Dataset as their corresponding thresholds.
-We calculate the proportion $P$ of samples in the Training Dataset with uncertainty values lower than these thresholds.
-If $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ are probabilistically independent, this proportion should be 81%.
-For other models with only one UQ module, we employ the $P$-th percentile of the uncertainty distribution from the Training Dataset as a threshold.
-The table shows the proportion of samples in the Uncertainty Samples that exceed this threshold.
-The higher the proportion, the stronger the model’s ability to identify uncertain samples.
-For each benchmark in the table, we highlight the model with the most effective filtering in green and the second most effective in blue.
-Note that $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, as parts of the Hybrid, are excluded from this ranking.
+The uncertainty threshold is established following the method described in subsubsection IV-C4, with the parameter $\alpha$ set to $1.3$.
+Figure 8: Uncertainty distributions produced by various BE-UQ models for the TPC-H+ benchmark.
+V-B4 Hardware & Software
+The experiments were carried out on a Ubuntu server with two Intel Xeon 4210R CPUs (10 cores, 2.40GHz), 256GB of main memory and one NVIDIA GeForce RTX 3060 GPU. The DBMS is PostgreSQL 12.13.
+V-C Performance of Uncertainty Quantification
+TABLE III: Proportion of unreliable samples in $D_{eval}$ successfully identified by various UQ methods.
+LIB-based
+Model
+TPC-H+ TPC-DS+ JOB+
+AMA-based
+Model
+TPC-H+ TPC-DS+ JOB+
+BNN 12.08 5.41 1.41 BNN 3.39 71.13 38.17
+MCD-all 67.95 39.16 20.60 MCD-all 19.30 69.33 22.36
+MCD-last 60.95 11.79 1.67 MCD-last 10.51 14.73 7.31
+MCD-dpp 7.74 5.54 2.16 MCD-dpp 76.50 11.25 13.20
+Ensemble 75.30 60.77 24.86 Ensemble 7.72 50.94 53.10
+AE 76.21 62.71 24.33 AE 77.01 76.90 53.75
+Hybrid 83.69 74.38 30.68 Hybrid 91.40 89.37 59.81
+STO 68.81 48.81 18.21
+HSTO 62.50 39.33 19.08
+In this subsection, we evaluate the uncertainty quantification (UQ) performance of various BE-UQ models.
+Evaluation Metric.
+We infer all BE-UQ models on $D_{train}$, $D_{test}$, and $D_{eval}$, obtaining the BE error and uncertainty of each sample in these datasets.
+We establish the BE error threshold for determining the reliability of a model’s BE results as the 90th percentile of BE errors on $D_{train}$.
+Samples in $D_{eval}$ with BE errors exceeding this threshold are designated as unreliable samples.
+To quantitatively evaluate the UQ performance of a BE-UQ model, we use the proportion of unreliable samples that are identified as unreliable estimates, that is, with uncertainties exceeding the specified uncertainty threshold, as the evaluation metric.
+A higher proportion indicates a better recall of unreliable samples realized by the BE-UQ model, thereby a higher UQ capability of the model.
+Experimental Results.
+The experimental results are presented in Table III.
+For each benchmark, the most effective BE-UQ model is highlighted in green, and the second most effective one is highlighted in blue.
+Furthermore, Figure 8 shows the uncertainty distributions produced by various BE-UQ models on $D_{train}$, $D_{test}$, and $D_{eval}$ using Raincloud plots.
+In each Raincloud plot, the horizontal axis is the uncertainty value, and the red line connects the means of uncertainty distributions.
+Due to space constraints, we only illustrate the results for the TPC-H+ benchmark; the results for other benchmarks are available in our open-source repository.
+From these results, we derive the following conclusions:
+(1) The Hybrid and AE models consistently outperform other methods, followed by the Ensemble approach.
+The remaining methods exhibit significantly inferior performance in both stability and unreliable sample detection capabilities.
+The Hybrid method achieves superior results across all benchmarks, particularly within the TPC-H+ benchmark, where the AMA-Hybrid model successfully identifies more than $91\%$ of unreliable samples, substantially surpassing alternative approaches.
+In more than five-sixths of evaluation scenarios, the AE framework demonstrates performance exceeding that of the Ensemble method, exhibiting superior stability.
+HSTO, despite being specifically engineered for Self-Attention mechanisms, fails to demonstrate enhanced effectiveness in this particular task.
+Among the three MCD-based methodologies, MCD-all demonstrates superior efficacy, achieving optimal results in five of the six evaluation scenarios.
+(2) Figure 8 shows that the uncertainties of the samples in $D_{train}$ mostly follow long-tailed distributions, which is consistent with the claim in subsubsection IV-C4 that the uncertainties follow long-tailed distributions.
+This verifies the rationale behind our method for setting uncertain thresholds.
+(3) The quality of UQ is fundamentally related to both the architecture of the base model and the benchmark.
+Across all three benchmarks, the UQ quality of the AMA base model consistently surpasses that of the LIB model, which is particularly evident in the proposed AE and Hybrid models.
+This phenomenon can potentially be attributed to the fact that a model’s capacity to effectively fit the data inherently influences both its accuracy and UQ capability.
+As demonstrated in Table IV, AMA’s accuracy significantly exceeds that of LIB.
+Accurate predictions necessitate effective encoding of inputs into a hidden vector representation, which constitutes the foundational principle of our methodology and accounts for these observed performance differentials.
+In summary, these $16$ models demonstrate superior performance on the TPC-H+ benchmark compared to TPC-DS+, with the least effective performance observed on the JOB+ benchmark, where even the most effective UQ methods filter only approximately $60\%$ of unreliable samples.
+TABLE IV: Percentiles of BE errors produced by various BE-UQ models when using both BE-UQ models and what-if tools for BE.The best-performing models are marked in green.
+Dataset LIB BNN MCD-all MCD-last MCD-dpp Ensemble STO HSTO AE Hybrid AMA BNN MCD-all MCD-last MCD-dpp Ensemble AE Hybrid What-if
+TPC-H+ 50th 0.41 0.56 0.12 0.19 0.36 0.09 0.34 0.18 0.09 0.07 0.28 0.31 0.33 0.32 0.08 0.36 0.05 0.04 0.05
+95th 0.92 0.77 0.97 0.97 0.94 0.77 0.68 0.91 0.68 0.59 0.74 0.72 0.73 0.76 0.68 0.73 0.53 0.42 0.60
+99th 0.98 1.03 0.82 0.97 0.95 0.89 0.81 0.95 0.83 0.79 0.86 0.78 0.80 0.90 0.78 1.07 0.62 0.62 0.78
+TPC-DS+ 50th 0.34 0.29 0.20 0.26 0.28 0.18 0.25 0.26 0.24 0.15 0.20 0.10 0.10 0.22 0.19 0.11 0.11 0.12 0.16
+95th 0.93 0.92 0.94 0.97 0.94 0.93 0.95 0.94 0.85 0.94 0.97 0.96 0.95 0.95 0.96 0.95 0.88 0.83 0.91
+99th 1.14 1.16 1.40 1.14 1.28 1.42 0.95 1.22 1.09 1.12 1.16 1.19 1.39 1.13 1.33 1.44 1.11 1.04 1.62
+JOB+ 50th 0.31 0.39 0.27 0.33 0.33 0.26 0.29 0.31 0.21 0.20 0.20 0.14 0.16 0.17 0.22 0.19 0.10 0.13 0.11
+95th 0.84 0.69 0.89 0.91 0.94 0.77 0.87 0.83 0.88 0.79 0.95 0.71 0.83 0.70 0.75 0.71 0.73 0.68 0.82
+99th 2.07 2.08 2.02 2.21 2.14 2.08 2.09 2.19 2.05 2.17 2.52 2.07 2.25 2.23 1.91 1.94 1.78 1.85 2.20
+TABLE V: Model inference time per batch (ms).
+Model TPC-H+ TPC-DS+ JOB+ Model TPC-H+ TPC-DS+ JOB+
+LIB 2.76 6.00 5.58 AMA 0.18 0.54 0.25
+BNN 61.85 156.73 50.75 BNN 10.77 20.01 10.54
+MCD-all 184.59 96.80 155.14 MCD-all 6.19 12.09 4.90
+MCD-last 60.00 94.72 133.04 MCD-last 5.23 11.68 5.23
+MCD-dpp 60.01 121.35 83.63 MCD-dpp 260.82 183.33 51.45
+Ensemble 22.72 32.48 29.90 Ensemble 1.82 3.14 1.71
+AE 3.57 5.11 6.37 AE 0.52 0.87 0.51
+Hybrid 60.74 81.32 98.16 Hybrid 10.28 16.45 9.26
+STO 151.40 185.76 157.01
+HSTO 64.46 167.11 327.03
+TABLE VI: Normalized benefit estimation time of various BE-UQ models. The estimation time using only what-if tools is normalized to 100% for each benchmark.
+Model TPC-H+ TPC-DS+ JOB+ Model TPC-H+ TPC-DS+ JOB+
+LIB 1.51% 0.72% 0.45% AMA 100.16% 100.07% 100.01%
+BNN 42.87% 36.99% 73.40% BNN 109.77% 102.41% 100.46%
+MCD-all 118.38% 34.77% 17.96% MCD-all 105.56% 101.45% 100.21%
+MCD-last 48.16% 33.41% 12.39% MCD-last 104.74% 101.40% 100.23%
+MCD-dpp 41.51% 34.52% 10.80% MCD-dpp 336.81% 122.10% 102.27%
+Ensemble 28.39% 30.91% 14.72% Ensemble 101.65% 100.37% 100.07%
+AE 27.77% 23.00% 3.68% AE 100.05% 100.10% 100.02%
+Hybrid 53.30% 33.70% 23.54% Hybrid 109.33% 101.98% 100.41%
+STO 98.86% 47.49% 21.15%
+HSTO 50.71% 43.04% 27.78%
+V-D Performance of Benefit Estimation
+In this subsection, we evaluate the benefit estimation (BE) performance of various BE-UQ models.
+Evaluation Metric. Given a sample $(q,\mathcal{I})$ as input, the BE error produced by a BE-UQ model is
+$\frac{|\hat{B}(q,\mathcal{I}_{0},\mathcal{I})-B(q,\mathcal{I}_{0},\mathcal{I})|}{c(q,\mathcal{I}_{0})},$
+where $\hat{B}(q,\mathcal{I}_{0},\mathcal{I})$ is the predicted benefit, $B(q,\mathcal{I}_{0},\mathcal{I})$ is the actual benefit, and $c(q,\mathcal{I}_{0})$ is the current cost of $q$.
+We evaluate the BE performance of a BE-UQ model by the 50th, 95th and 99th percentiles of the BE errors produced by the BE-UQ model for all samples in $D_{eval}$.
+We do not use the average BE error as the evaluation metric because a few very large errors do not substantially affect the average error but significantly degrade the performance of index tuning.
+Experimental Results.
+Table IV presents the percentiles of BE errors generated by various BE methodologies, encompassing what-if tools, learning-based BE models, and BE-UQ models.
+BE-UQ models augment their predictions with results from what-if tools when unreliable samples are identified, as illustrated in Figure 6.
+This experiment directly reflects accuracy when BE-UQ models are integrated into the index tuning process.
+For each benchmark, the best-performing paradigm is highlighted in green.
+Models derived from different base models are evaluated independently.
+In conjunction with Table III, two principal findings emerge:
+(1) Precise uncertainty quantification substantially enhances the accuracy of benefit estimation.
+As demonstrated in subsection V-C, the proposed Hybrid and AE methods exhibited superior UQ performance.
+Consequently, in Table IV, the Hybrid approach outperforms learning-based BE models in $89\%$ of metrics, exceeds what-if tools in $72\%$ of metrics, and surpasses all other BE-UQ methods in $61\%$ of metrics.
+Models employing the AE technique demonstrate marginally lower accuracy than the Hybrid approach but consistently outperform other UQ methods.
+(2) Deficient UQ capabilities in BE-UQ architectures combine the limitations of both learning-based BE models and what-if tools, thereby compromising the accuracy of benefit estimation.
+A notable example is MCD-last, which, as indicated in Table III, filters out merely $1.67\%$ of unreliable results during its worst performance.
+This leads to a disproportionate elimination of accurate estimation results during benefit estimation, while inaccurate results provided by the what-if tools are retained, culminating in suboptimal performance.
+Our analyses demonstrate that the impact of UQ on BE accuracy is fundamentally dependent upon the robustness of UQ capabilities.
+When UQ capabilities are robust, BE-UQ models effectively synthesize the strengths of learning-based models and what-if tools, yielding superior results compared to either approach in isolation.
+Conversely, when UQ capabilities are deficient, BE-UQ accuracy typically falls between these two methods, and in certain scenarios, may underperform both approaches.
+V-E Performance of Inference Efficiency
+In this subsection, we evaluate the inference efficiency of various BE-UQ models.
+Experimental Results (Using BE-UQ Models Only).
+We evaluate computational efficiency of various models by measuring the time to process a batch of samples, encompassing both benefit estimation and uncertainty quantification phases.
+The experimental results are presented in Table V.
+The BE-UQ models that achieve efficiency comparable to the base models within the same order of magnitude are highlighted in green.
+(1) It is evident that among all UQ methods, AE demonstrates the highest efficiency, followed by Ensemble.
+Compared to the base model, AE necessitates only one additional inference for the decoder network.
+In contrast, the inference time for Ensemble increases linearly with the number of models.
+Nevertheless, Ensemble methods still demonstrate higher efficiency compared to other multiple-inference UQ techniques, such as BNN, MCD, HSTO, and Hybrid.
+(2) Model inference efficiency is influenced by the base model structure and hyperparameters.
+For instance, LIB, which uses a heavier Multi-head Attention mechanism, requires up to an order of magnitude more inference time than AMA.
+For models based on the same base architecture, the optimal hyperparameters differ across various benchmarks, leading to notable differences in inference efficiency.
+Experimental Results (Using Both BE-UQ Models and What-if Tools).
+We evaluated various models under the Beauty framework for BE with an equivalent number of samples.
+We recorded the BE call sequence during an index tuning process and measured the time required by what-if tools and other models to complete the BE sequence, as illustrated in Table VI.
+For each benchmark, we established what-if tools as the baseline, setting their estimation time to $100\%$.
+This assessment reflects the practical efficiency of the models when applied to index tuning tasks.
+The BE-UQ model with the shortest inference time is highlighted in green.
+From experimental results, it is concluded that, after using the Beauty framework, the primary factor influencing BE efficiency is not the model’s inference time, but rather the number of what-if calls.
+This is due to the substantially lower efficiency of the what-if tools compared to learning-based models.
+Specifically, each what-if call requires tens to hundreds of milliseconds [1], and unlike GPU-accelerated batch processing in learning-based models, the what-if tools execute sequentially without hardware acceleration.
+(1) For LIB-based models, only unreliable results from models require supplementation by what-if tools, which leads to a lower inference time compared to the what-if tools baseline.
+The AE remains the most efficient, but its performance improves from being one to two orders of magnitude faster than other methods to being several times faster.
+(2) In contrast, models based on AMA necessitate query plans generated by what-if tools as input to obtain BE results.
+Consequently, when compared to baselines that utilize what-if tools, AMA-based models demonstrate marginally slower BE times in index tuning scenarios.
+Upon implementation of the Beauty framework, BE-UQ models incur additional computational overhead relative to baseline, primarily comprising network inference latency and UQ processing time.
+With the exception of MCD-dpp, which requires diversity sampling, the computational cost of the UQ for other AMA-based models remains negligible.
+V-F Section Conclusion
+Synthesizing the findings presented in Tables III through VI, we can draw the following conclusions: With respect to UQ capabilities, the Hybrid demonstrates superior performance, followed by the AE, both of which surpass alternative methods.
+This enhanced UQ capability translates into significant improvements in BE accuracy when integrated into the Beauty framework.
+In terms of efficiency, AE emerges as the frontrunner, outperforming both the Hybrid and other methods.
+Consequently, the optimal choice between Hybrid and AE is contingent upon the specific application scenario.
+In scenarios where time-constrained tuning is paramount [35], AE represents the preferred option.
+For all other scenarios, the Hybrid approach is recommended as the primary selection.
+VI Empirical Study on Index Tuning
+In this section, we compare the performance of what-if tools, learning-based BE models, and BE-UQ models in index tuning, to investigate whether UQ can enhance the results of index tuning.
+VI-A Experimental Setup
+VI-A1 Workloads
+In addition to TPC-H+, TPC-DS+, and JOB+, we also conduct index tuning on workloads containing only template-generated queries without synthetic queries, namely TPC-H, TPC-DS, and JOB.
+VI-A2 Implementation
+Leis et al.
+[32] demonstrate that no single index configuration enumeration algorithm delivers optimal performance across all scenarios.
+In our experiments, we employed DTA [13], which effectively balances runtime and solution quality across various contexts.
+This algorithm operates with a predefined maximum runtime and terminates after enumerating all configurations or reaching this time constraint.
+However, following Leis et al.’s [32] implementation, the algorithm completes the current seed’s configuration even when exceeding the time limit, occasionally surpassing the maximum allocated runtime.
+We calibrate the maximum runtime according to workload size, as detailed in Table II.
+Figure 9: Cost improvements achieved by different estimators.
+TABLE VII: Rankings of cost improvements achieved by different benefit estimators.
+Method # 1st Place # 2nd Place # 3rd Place # Worst
+LIB 1 1 0 21
+LIB-Ensemble 2 1 5 5
+LIB-AE 2 1 5 2
+LIB-Hybrid 2 4 8 3
+AMA 4 2 3 20
+AMA-Ensemble 3 7 1 13
+AMA-AE 8 8 8 7
+AMA-Hybrid 15 9 6 0
+What-if 5 9 8 2
+TABLE VIII: Statistical information on cost improvement of different benefit estimators across various benchmarks.
+Method TPC-H TPC-DS JOB TPC-H+ TPC-DS+ JOB+
+Mean Var Mean Var Mean Var Mean Var Mean Var Mean Var
+LIB 0.029 $7.9\cdot 10^{-4}$ 0.079 0.011 0.352 0.093 $9.5\cdot 10^{-4}$ $7.0\cdot 10^{-5}$ 0.040 0.013 0.370 0.020
+LIB-Ensemble 0.041 0.002 0.216 0.012 0.555 0.019 0.001 $1.0\cdot 10^{-5}$ 0.194 0.009 0.586 0.009
+LIB-AE 0.033 $3.0\cdot 10^{-4}$ 0.242 0.004 0.670 0.008 0.004 $2.4\cdot 10^{-5}$ 0.266 0.001 0.437 0.027
+LIB-Hybrid 0.082 0.006 0.163 0.020 0.535 0.061 0.039 0.002 0.222 0.003 0.554 0.012
+AMA 0.009 $3.3\cdot 10^{-4}$ 0.153 0.029 0.535 0.116 0.017 0.001 0.120 0.017 0.226 0.084
+AMA-Ensemble 0.011 $2.4\cdot 10^{-4}$ 0.255 0.012 0.586 0.109 $0.009$ $1.6\cdot 10^{-4}$ 0.164 0.018 0.575 0.008
+AMA-AE 0.085 0.005 0.258 0.011 0.542 0.185 0.059 0.002 0.287 0.002 0.602 0.002
+AMA-Hybrid 0.105 0.006 0.306 0.004 0.688 0.042 0.075 0.003 0.310 0.009 0.542 0.013
+What-if 0.083 0.004 0.250 0.011 0.687 0.046 0.049 $2.8\cdot 10^{-4}$ 0.251 0.009 0.493 0.078
+Based on the performance of various models discussed in Section V, we selected nine models for this experiment: LIB, LIB-Ensemble, LIB-AE, LIB-Hybrid, AMA, AMA-Ensemble, AMA-AE, AMA-Hybrid, and what-if.
+The thresholds for each model were established as described in Subsection V-B3.
+When the uncertainty value of benefits estimated by the BE-UQ model for a sample exceeds the threshold, we utilize what-if tools to estimate the benefits for that particular sample.
+VI-A3 Evaluation Metrics
+Given a workload $W$, $\mathcal{I}_{0}$ represents the initial set of indexes, and $\mathcal{I}$ denotes the set of selected indexes.
+We quantify the quality of replacing $\mathcal{I}_{0}$ with $\mathcal{I}$ by the relative improvement in $W$’s cost: $\max\left(0,\frac{B(W,\mathcal{I}_{0},\mathcal{I})}{c(W,\mathcal{I}_{0})}\right)=\max\left(0,1-\frac{c(W,\mathcal{I})}{c(W,\mathcal{I}_{0})}\right).$
+In scenarios where the selected index configuration leads to an increase in the total workload cost, such configurations must be discarded.
+Therefore, in such cases, the improvement in Figure 9 will be marked as zero.
+If multiple estimators have zero improvement under the same budget, the “# Worst” in Table VII will count them repeatedly.
+VI-B Results on Index Tuning
+Figure 9 illustrates the index tuning results of nine methods under various budgets across six benchmarks.
+Table VII presents the ranking information of each benefit estimator across the six benchmarks, highlighting the best performances in green.
+Furthermore, Table VIII demonstrates the mean and variance of cost improvement achieved by each benefit estimator across every benchmark.
 From these results, we can draw the following conclusions:
-First, the AE and Hybrid models consistently perform the best, followed by the Ensemble method.
-The other methods lag significantly behind in both stability and optimal performance in UQ.
-The best performance is seen with the AMA-AE model under the TPC-H+, which can identify 92.68% of Uncertainty Samples while misclassifying about 19% of training samples, significantly outperforming general-purpose methods.
-In more than two-thirds of the benchmarks, the AE and Hybrid models consistently rank among the top two, demonstrating better stability than the Ensemble method.
-Second, the quality of uncertainty quantification is related to the base model and dataset.
-Across three benchmarks, the uncertainty modeling quality of the AMA base model consistently surpasses that of the LIB model, particularly evident in the proposed AE and Hybrid models.
-This may be attributed to the fact that the model’s ability to fit the data also determines its accuracy and capability in UQ.
-As illustrated in Table V, AMA’s accuracy is significantly higher than LIB’s.
-Accurate results necessitate effective encoding of inputs into a hidden vector, the foundational principle of our methodology, which accounts for these observations.
-Overall, these 16 models demonstrate superior performance on the TPC-H+ bencamark compared to TPC-DS+, with the poorest performance observed on the JOB+ benchmark, where the most effective UQ methods filter only 61% of Uncertainty Samples.
-Third, the settings for uncertainty thresholds significantly impact the ability to detect Uncertainty Samples.
-Uncertainty distribution within a dataset is usually confined to a narrow range, and the relationship between the threshold and this range dictates the model’s detection capabilities.
-For example, in the AMA-Ensemble model under TPC-H+, setting a filtering threshold at the 81th percentile of the training dataset excludes only 38.73% of Uncertainty Samples, whereas a threshold at the 75th percentile increases exclusion to 73.34%.
-Fourth, HSTO, specifically designed for Self-Attention, did not demonstrate improved performance on this task. Neither model achieved optimal performance across the three benchmarks.
-Fifth, among the three MCD-based methods, MCD-all outperforms the others, achieving optimal results in five out of six scenarios.
-MCD-dpp provides a richer sampling diversity compared to MCD-last and outperforms it in the majority of scenarios.
-TABLE IV: Time for model inference per batch (ms).
-Model TPC-H+ TPC-DS+ JOB+ Model TPC-H+ TPC-DS+ JOB+
-LIB 2.76 2.30 5.58 AMA 0.18 0.18 0.25
-BNN 61.85 44.46 50.75 BNN 10.77 7.41 10.54
-MCD-all 184.59 45.03 155.14 MCD-all 6.19 2.94 4.90
-MCD-last 60.00 89.63 133.04 MCD-last 5.23 4.13 5.23
-MCD-dpp 60.01 147.86 83.63 MCD-dpp 260.82 78.39 51.45
-Ensemble 22.72 14.15 29.90 Ensemble 1.82 0.80 1.71
-AE 3.08 4.76 6.37 AE 0.52 0.51 0.51
-Hybrid 60.74 103.95 98.16 Hybrid 10.28 9.82 9.26
-STO 151.40 137.33 157.01
-HSTO 64.46 109.25 327.03
-TABLE V: The impact of uncertainty quantification methods on the prediction errors.
-Models that have the most negative impact are marked in red, while those with the most positive impact are marked in green.
-Dataset LIB BNN MCD-all MCD-last MCD-dpp Ensemble STO HSTO AE Hybrid AMA BNN MCD-all MCD-last MCD-dpp Ensemble AE Hybrid
-TPC-H+ 50th 0.13 +0.10 +0.01 +0.02 +0.04 +0.00 +0.03 +0.00 +0.09 +0.05 0.01 +0.02 +0.00 +0.00 +0.02 +0.02 +0.00 +0.00
-95th 0.47 +0.02 +0.00 +0.00 -0.04 +0.00 +0.00 +0.00 +0.05 +0.01 0.05 +0.06 +0.01 +0.03 +0.00 +0.03 +0.00 +0.02
-99th 0.65 -0.11 -0.05 -0.04 -0.07 -0.01 +0.00 -0.05 -0.12 -0.01 0.11 +0.07 +0.00 +0.00 +0.03 +0.01 -0.03 +0.02
-TPC-DS+ 50th 0.06 +0.01 +0.00 -0.01 +0.00 +0.00 +0.00 +0.00 +0.00 +0.00 0.01 +0.01 +0.00 +0.00 +0.00 +0.00 +0.00 +0.00
-95th 0.44 +0.03 +0.00 -0.02 +0.01 +0.00 +0.00 -0.02 +0.00 -0.04 0.20 +0.06 +0.02 +0.00 +0.09 +0.03 +0.01 +0.01
-99th 0.71 +0.00 +0.00 +0.03 +0.00 +0.00 +0.00 +0.00 +0.02 +0.00 0.34 +0.08 +0.06 +0.05 +0.07 +0.04 +0.05 +0.07
-JOB+ 50th 0.19 +0.13 +0.01 +0.03 +0.04 +0.00 +0.04 +0.05 +0.03 +0.01 0.03 +0.02 +0.00 +0.00 +0.05 +0.00 +0.00 +0.01
-95th 0.57 +0.04 +0.00 +0.00 +0.00 +0.00 +0.02 +0.00 +0.03 +0.00 0.19 +0.06 -0.01 +0.00 +0.05 +0.00 +0.00 +0.04
-99th 0.76 +0.16 -0.02 +0.02 +0.00 -0.02 +0.07 +0.09 +0.04 +0.04 0.37 +0.17 +0.00 +0.00 +0.11 +0.00 +0.08 +0.10
-5.4 Results on Prediction Accuracy and Inference Efficiency
-Table IV shows the time taken by different models to process a single batch, which includes predicting the relative error $\hat{rb}_{predicted}$ and estimating uncertainty, in milliseconds.
-Models in which the inference time with UQ is on the same order of magnitude as the base model are marked in green.
-The primary factors influencing inference efficiency are the base model structure, model hyperparameters, and methods for UQ.
-LIB, utilizing heavier Multi-head Attention, requires significantly more inference time, up to an order of magnitude higher than AMA.For the models based on the same BE model, the optimal hyperparameters vary across different datasets, which can also lead to several times difference in inference efficiency.
-Methods like BNN, MCD, and HSTO necessitate multiple inferences on the same model, resulting in inference times that are proportionally related to the number of inferences, significantly exceeding the base model and AE by one to two orders of magnitude.
-Specifically, MCD-dpp introduces extra computational overhead due to sampling processes, notably affecting the faster base model AMA, and to a lesser extent, the heavier LIB model where it is less of a determining efficiency factor.
-The inference time for Ensemble models scales with the number of models involved; selecting five models during hyperparameter tuning has resulted in higher overall efficiency compared to other multiple-inference UQ methods.
-Of all methods evaluated, AE proves the most efficient, necessitating only an additional inference for the decoder network, which keeps its inference times close to those of the base model on all six benchmarks.
-Table V evaluates the impact of incorporating UQ methods into BE models.
-We assess performance using the absolute difference between predicted and actual values, i.e.
-$|rb_{actual}-\hat{rb}_{predicted}|$.
-The table shows the 50th, 95th, and 99th percentiles of error distributions for LIB and AMA.
-For models to which UQ methods have been added, the table shows the change in error relative to the base model, where models that have the most negative impact are marked in red, and those with the most positive impact are marked in green.
-Among all UQ methods, only BNN significantly affected the accuracy of the BE model, while the other methods had no substantial impact.
-Each model was retrained using 30 distinct sets of hyperparameters, so any minor variations in accuracy can be attributed to differences in hyperparameter settings.
-In the case of BNN, however, each neuron is replaced by a value sampled from a distribution, which effectively doubles the number of network parameters (as each deterministic neuron corresponds to values such as mean, variance, etc., in BNN) and significantly expands the hyperparameter space (e.g., requiring a prior distribution).
-This makes the training process fundamentally different and more complex, requiring more intensive hyperparameter tuning.
-Comparing to the MCD-last, MCD-dpp had a higher effect on accuracy when AMA was used as the base model, more inference burden while more UQ capability, making it better suited for integration with heavier base models.
-, which might be related to the changes it made in the sampling process during inference.
-However, more research is needed to provide a definitive conclusion, which is beyond the scope of this study.
-Among the other methods, since the AE method did not alter the original structure and parameters of the BE model after training, it theoretically should not affect the accuracy of the BE model.
-6 Empirical Study on Index Tuning
-6.1 Experimental Setup
-In this section, we explore whether modeling uncertainty can enhance index tuning results.
-Therefore, we integrated the aforementioned BE model into the Beauty framework and performed index tuning on six benchmarks under different budget constraints.
-6.1.1 Implementation
-Leis et al.
-[38] show that no single index configuration enumeration algorithm delivers optimal performance in all scenarios.
-In our experiments, we used DTA [14], which balances runtime and solution quality well across various scenarios.
-This algorithm has a set maximum runtime and stops after enumerating all configurations or reaching this limit.
-However, following Leis et al.’s [38] implementation, the algorithm completes the current seed’s configuration even if it exceeds the time limit, occasionally surpassing the maximum runtime.
-We adjust the maximum runtime based on workload size, as shown in Table II.
-Based on the performance of various models mentioned in Section 5, we selected nine models in our experiment: LIB, LIB-Ensemble, LIB-AE, LIB-Hybrid, AMA, AMA-Ensemble, AMA-AE, AMA-Hybrid, and what-if.
-The thresholds for each model were set as described in Subsection 5.3.
-When the uncertainty value of the benefits estimated by the BE-UQ model for a sample exceeds the threshold, we use what-if tools to estimate the benefits of that sample.
-6.1.2 Evaluation Metrics
-We use specific metrics to assess the performance of index tuning.
-Given a workload $W$, $\mathcal{I}_{0}$ represents the initial set of indexes, and $\mathcal{I}$ the set of selected indexes.
-We measure the quality of replacing $\mathcal{I}_{0}$ with $\mathcal{I}$ by the relative improvement in $W$’s cost: $\max\left(0,\frac{B(W,\mathcal{I}_{0},\mathcal{I})}{c(W,\mathcal{I}_{0})}\right)=\max\left(0,1-\frac{c(W,\mathcal{I})}{c(W,\mathcal{I}_{0})}\right).$
-On some occasions, if the selected indexes increase the cost, they must be discarded, and thus the improvement in cost is zero.
-Figure 11: Cost improvements achieved by different benefit estimation models.
-TABLE VI: Rankings of cost improvements achieved by models.
-Method # 1st Place # 2nd Place # 3rd Place # Worst
-What-if 8 8 11 1
-LIB 0 0 1 10
-LIB-Ensemble 1 4 4 3
-LIB-AE 3 1 1 4
-LIB-Hybrid 3 7 0 4
-AMA 3 3 3 6
-AMA-Ensemble 4 2 6 7
-AMA-AE 10 7 5 0
-AMA-Hybrid 4 4 5 0
-6.2 Results on Index Tuning
-Figure 11 illustrates the index tuning results of nine methods under various budgets across six benchmarks.
-Table VI presents detailed statistics of cost improvements, highlighting the best performances in green and the second best in blue.
-From these results, we can draw the following conclusions:
-First, integrating the UQ module into the two base models significantly improved index tuning results.
-The LIB model has seen a significant reduction in the number of worst-case scenarios, while the AMA model has experienced a substantial increase in the number of optimal configurations.
-Additionally, when integrating AE and Hybrid into the AMA model, the number of worst-case scenarios decreased from six to zero.
-Second, the performance of the integrated Hybrid model is superior to that of the Ensemble-based model.
-As shown in Table III, Hybrid’s ability to quantify uncertainty surpasses that of Ensemble, allowing it to more accurately identify the parts of the learning-based BE model that it cannot accurately estimate, thereby improving BE’s accuracy.
-The DTA algorithm requires a specified maximum runtime, and usually, it cannot enumerate all index configurations by the end.
-Therefore, the greater the delay in BE-UQ inference, the fewer configurations the index advisor can explore.
-As seen in Table IV, Hybrid has a higher delay.
-Thus, by providing more accurate uncertainty quantifications, Hybrid finds better results with fewer configurations.
-Third, AE’s results are slightly better than Hybrid’s.
-The accuracy of the LIB base model is not high, and in this situation, Hybrid’s UQ accuracy is higher than AE’s, enabling it to identify as many uncertain data points as possible, resulting in better performance than AE.
-Conversely, when using AMA as the base model, AE achieves higher identification accuracy and lower inference delay, allowing for more accurate enumeration of more index configurations.
-Therefore, it achieves the highest number of optimal configurations and surpasses the what-if tools.
-For index tuning tasks, Hybrid is preferable when there is sufficient time available for tuning.
-If time is limited and the BE model demonstrates high accuracy, selecting AE can result in a better index configuration.
-Fourth, in scenarios abundant with Uncertainty Samples, the what-if approach remains a stable and precise method, achieving the best or nearly the best results across all six benchmarks.
-This consistency forms the foundation of our study, which aims to model the uncertainty in learning-based models, combining their accuracy and efficiency with the reliability of what-if tools.
-7 Related Work
-We presented the related work on UQ in subsection 2.3 and subsection 5.2.
-The related work on learning-based BE models is detailed in subsection 2.2.
-In this section, we introduce additional related work on index tuning.
-Index Configuration Enumeration.
+Firstly, all three UQ methods enhance index tuning results compared to the base model: the frequency of worst-case scenarios decreases markedly, while the occurrence of optimal configurations increases.
+As evident from Table VIII, applying any UQ method to the base model yields improved cost performance, particularly with Hybrid and AE methodologies.
+This addresses the question: uncertainty quantification indeed enhances learning-based benefit estimation, resulting in superior index tuning outcomes.
+Secondly, models incorporating the Hybrid approach demonstrate superior performance.
+Notably, AMA-Hybrid substantially outperforms other models in both cost improvement rankings and mean values.
+As illustrated in Table III, Hybrid’s capability to quantify uncertainty exceeds that of Ensemble and AE methods, enabling more precise identification of samples within the learning-based BE framework that cannot be accurately estimated, thereby enhancing BE’s precision.
+The DTA algorithm necessitates a predefined maximum runtime, and typically cannot enumerate all index configurations before termination.
+Consequently, greater delays in BE-UQ inference restrict the number of configurations the index advisor can explore.
+As evidenced in Table V, Hybrid exhibits increased latency.
+Nevertheless, by providing more accurate uncertainty quantifications, Hybrid achieves superior results with fewer configuration evaluations.
+Thirdly, AE’s results surpass those of Ensemble.
+According to Table VIII, in $75\%$ of cases, the mean cost improvement when integrating AE exceeds that of Ensemble.
+This is further reflected in Table VII, wherein AE achieves significantly better rankings than Ensemble.
+These findings validate the effectiveness of our proposed UQ method based on the AutoEncoder architecture.
+Fourthly, in scenarios replete with unreliable results, the what-if tools remain a stable and accurate methodology, significantly outperforming both base BE models.
+The effectiveness of these base BE models exhibits heightened sensitivity to benchmarks and storage constraints, thereby demonstrating insufficient robustness.
+This consistency forms the foundation of our investigation, which aims to model uncertainty in learning-based models, combining their computational efficiency and predictive accuracy with the reliability inherent in what-if tools.
+VII Related Work
+We presented the related work on UQ in subsection II-C and subsection V-B.
+The related work on learning-based BE models is detailed in subsection II-B.
+This section focuses on the works concerning configuration enumeration algorithms.
 Early greedy enumeration algorithms typically follow two strategies [32].
@@
 SmartIX [37] introduces add/delete operations based on environmental state.
-Workload Compression.
-To enhance the efficiency of index tuning, many approaches perform workload compression prior to tuning.
-Sadri et al.
-[41] identify templates based on the non-parameterized portions of query statements, retaining one query for each template.
-However, the query plans of a template can vary greatly depending on parameter differences, resulting in substantially different benefits calculated for the same index configuration and causing significant workload information loss.
-Furthermore, RIBE [39] clusters queries based on operators within the query plan, ensuring that only redundant information from the workload is eliminated.
-Besides these clustering-based methods, GSUM [2] and ISUM [43] define various query features and design greedy algorithms that flexibly determine the number of queries in the workload.
-8 Conclusion
-In this paper, we address a critical yet unexplored question: Can uncertainty enable better learning-based index tuning?
-To answer this question, we focus on the core component of index tuning, prioritizing improvements to the learning-based BE model.
-Our investigation is divided into two sub-questions: (1) Can UQ methods accurately measure the reliability of model results?
-(2) How can a BE model that provides uncertainty estimates be integrated into an index advisor, and what are the implications of this integration?
-To address these two questions, we first analyze the origins and characteristics of model uncertainty in BE tasks.
-Unlike existing general UQ methods, we design a new hybrid quantification framework tailored to the task’s characteristics and propose an alternative version to enhance efficiency.
-Experimental validation shows that our method outperforms existing quantification methods in five out of six scenarios.
-The alternative version significantly improves UQ efficiency, albeit with a slight reduction in quantification accuracy.
-Next, we verified that the Beauty framework leverages the stability of what-if tools to supplement the learning-based BE model.
-In end-to-end experiments, this framework significantly improves the quality of index tuning results.
-Our findings confirm that integrating uncertainty can enhance learning-based index tuning.
-Our framework focuses on the BE problem, marking an initial exploration in this area.
-Future research will investigate the role of uncertainty in other index advisor modules.
+VIII Conclusion
+In this paper, we address a critical yet unexplored question: Can uncertainty quantification enable better learning-based index tuning?
+To answer this question, we focus on the core component of index tuning by prioritizing improvements to the learning-based BE model.
+Our investigation revolves around two sub-questions: (1) Can UQ methods accurately measure the reliability of model results?
+(2) How can a BE-UQ model be integrated into an index advisor?
+To address these questions, we first analyze the origins and characteristics of model uncertainty in BE tasks.
+We then design a novel hybrid quantification framework and propose an efficient variant.
+Experiments show that our method outperforms existing methods in most cases.
+Additionally, we demonstrate that the Beauty framework significantly improves the quality of index tuning results.
+While our framework focuses on the BE problem, it represents an initial exploration in this area. Future research will investigate the role of uncertainty in other modules of the index advisor.
 Acknowledgments
-This work was partially supported by the National Natural Science Foundation of China (Grant No. 62072138).
+This work was supported by the National Natural Science Foundation of China (No. 62072138, No. 62472123).